### v1

# ECG-Byte: A Tokenizer for End-to-End Generative Electrocardiogram Language Modeling

Abstract
Large Language Models (LLMs) have shown remarkable adaptability across domains beyond text, specifically electrocardiograms (ECGs). More specifically, there is a growing body of work exploring the task of generating text from a multi-channeled ECG and corresponding textual prompt. Current approaches typically involve pretraining an ECG-specific encoder with a self-supervised learning (SSL) objective and using the features output by the pretrained encoder to finetune a LLM for natural language generation (NLG). However, these methods are limited by 1) inefficiency from two-stage training and 2) interpretability challenges with encoder-generated features. To address these limitations, we introduce ECG-Byte, an adapted byte pair encoding (BPE) tokenizer pipeline for autoregressive language modeling of ECGs. This approach compresses and encodes ECG signals into tokens, enabling end-to-end LLM training by combining ECG and text tokens directly, while being much more interpretable since the ECG tokens can be directly mapped back to the original signal. Using ECG-Byte, we achieve competitive performance in NLG tasks in only half the time and $\sim$48of the data required by two-stage approaches.
Data and Code Availability
This paper uses the ECG-Chat pretraining (Zhao et al., 2024) and ECG-QA datasets (Oh et al., 2023), which were both created from the MIMIC-IV ECG (Johnson et al., 2023) and PTB-XL datasets (Wagner et al., 2020). More details about the datasets are provided in Section 3. The ECG-Chat pretraining, ECG-QA, MIMIC-IV ECG, and PTB-XL datasets are all freely available on https://github.com/YubaoZhao/ECG-Chat, https://github.com/Jwoo5/ecg-qa, and https://physionet.org/ respectively. Lastly, we have released the code at the following link: https://github.com/willxxy/ECG-Byte.
Institutional Review Board (IRB)
All datasets used in this study are directly taken from the publicly available, de-identified MIMIC-IV ECG (Johnson et al., 2023) and PTB-XL (Wagner et al., 2020) datasets, thus not requiring IRB approval.
1 Introduction
Cardiovascular diseases (CVDs) are the leading cause of global mortality, with 17.9 million lives taken each year and increasing (Organization, 2024). Due to their readily available, noninvasive and information dense nature, 12-lead ECGs are first-line diagnostic tools for screening/evaluation of potential CVDs. However, accurate ECG analysis is limited in places where ECG expertise is not accessible, exacerbated by the decline and lack of available cardiac electrophysiologists especially in rural areas (Johnson, 2024).
Figure 1: Comparison of traditional methods and our method on ECG language modeling. Traditional methods comprises 2 training stages. The first stage aims to learn a good representation of the 12-lead ECG by training an ECG-specific encoder with a self-supervised learning objective via a combinatorial or individual usage of contrastive learning ($L_{CL}$) on the hidden states of the ECG $z_{e}$ and text $z_{t}$ or masked image modeling ($L_{MIM}$). The second stage applies the trained encoder to an ECG, which is input alongside a text prompt to a LLM for generation. In contrast, our method is end-to-end and can directly train a LLM for generation utilizing ECG-Byte.
The aforementioned facts calls attention to the need for accessible, accurate, and efficient automation of ECG analysis through deep learning. Deep learning has reached expert level performance in certain tasks for CVD detection using ECGs (Rajpurkar et al., 2017; Hannun et al., 2019; Qiu et al., 2023b). However, most previous works in this domain have succumbed to a crude classification of hard CVD labels (Choi et al., 2023; Nonaka and Seita, 2020; Martin et al., 2021; Strodthoff et al., 2021). A problem with this approach is that ECGs often do not exclusively fall into one diagnostic category, instead, there may be many soft labels annotated by expert physicians and the accumulation of these soft labels allow a more detailed, nuanced, and clinically useful interpretation of the ECG.
The recent onset of Large Language Models (LLMs) provides an opportunity to take a softer, generative, and consequently more flexible approach to ECG analysis. There are some recent works that have taken advantage of this approach to ECG analysis (Qiu et al., 2023a; Tang et al., 2024b; Wan et al., 2024; Fu et al., 2024; Zhao et al., 2024). A commonality among these works is that they treat multi-channel ECGs similarly to images; they first pretrain an encoder specifically for ECGs with some self-supervised learning (SSL) objective, then apply the learned features to finetune a LLM for natural language generation (NLG). However, we observed 2 limitations with this approach: 1) inefficiencies from two-stage training and 2) interpretability challenges with encoder generated features. Pretraining a good ECG specific encoder can be a significant computational burden due to large datasets, model size, and long training times. Additionally, the latent feature vector output by the ECG encoder cannot be mapped back to the signal, making interpretability difficult when utilizing this feature vector for downstream tasks.
In this study, we introduce ECG-Byte, an adapted byte pair encoding (BPE) (Gage, 1994) tokenizer designed for end-to-end training for generative ECG language modeling. Inspired by prior works demonstrating the effectiveness of creating discrete tokens from continuous values (Chen et al., 2022; Han et al., 2024), we leverage quantization to represent amplitude ranges as discrete symbols. Using discrete symbols, we obtain string representations of the ECGs to train ECG-Byte. We then apply ECG-Byte to directly finetune a LLM for conditional autoregressive NLG, where the text output is conditioned on the text prompt and tokenized signal. We found that our end-to-end training approach is competitive in conditional NLG with only half the time and $\sim$48of the data required by 2-stage pretraining approaches. Additionally, since the encoded ECG can be reverse transformed back to the original signal, we can interpret token-level attention-based visualizations, whereas it would be impossible to reverse the hidden latent feature vector outputs by a pretrained encoder.
Our contributions are summarized as the following:
  1. 1.
    We present ECG-Byte, an adapted BPE tokenizer for end-to-end training on autoregressive, conditional NLG.
  2. 2.
    We empirically show the efficiency of our method and present competitive performance compared to conventional 2-stage pretraining approaches, proposing a new paradigm for conditional NLG with ECGs.
  3. 3.
    We conduct an interpretability study on both ECG-Byte and the LLM by respectively observing how ECG-Byte is merging ECG signals and by utilizing attention visualizations to observe how the LLM is processing ECGs and text.
2 Related Works
2.1 Deep learning for ECGs
There has been a plethora of works utilizing deep learning for processing ECGs for classification (Rajpurkar et al., 2017; Hannun et al., 2019; Choi et al., 2023; Nonaka and Seita, 2020; Martin et al., 2021; Strodthoff et al., 2021). Most of these works utilize either convolutional neural networks (CNNs) (Rajpurkar et al., 2017) or transformers (Choi et al., 2023) and exhibit excellent performance at classification tasks. There have also been some efforts to frame classification as a retrieval task in order to recover cases similar to the given ECG.(Tang et al., 2024b; Qiu et al., 2023b). However, the retrieval approach may struggle with rare or unique ECG patterns that lack good matches in the existing database, potentially leading to missed or inaccurate diagnoses for unusual cases. Additionally, both classification and retrieval tasks may be crude formulations of processing ECGs, since ECGs typically exhibit many characteristics of overlapping CVDs.
2.2 Large Language Models for ECGs
Generative Large Language Models (LLMs) have given the opportunity to take a softer and more clinically similar approach in processing ECGs by generating physician-vetted clinical statements (Qiu et al., 2023a; Tang et al., 2024b; Wan et al., 2024; Fu et al., 2024; Zhao et al., 2024). The representation of ECG data has been largely considered in previous works which feed the raw signal into an encoder to obtain a latent representation of the ECG data, which then serves as input into a LLM. (Zhao et al., 2024; Wan et al., 2024; Tang et al., 2024b; Choi et al., 2023). In order to get good latent representations of the ECGs, an encoder is first pretrained on a self-supervised learning (SSL) objective (e.g., contrastive learning, masked language/image modeling). Although model performance in terms of label classification has been excellent when using these approaches, we want to be able to generate soft labels akin to clinical notation since ECGs often have overlapping and non-mutually exclusive descriptors. There have been some efforts in this direction (Qiu et al., 2023a; Tang et al., 2024b; Wan et al., 2024; Fu et al., 2024; Zhao et al., 2024), however, they suffer in efficiency (i.e., requires 2 stages of training). In our work, we challenge this 2-stage pretraining approach by transforming the ECG into tokens using ECG-Byte and directly training a LLM for NLG.
2.3 Byte Pair Encoding for Domains Outside of Language
The Byte Pair Encoding (BPE) algorithm was first introduced by Gage (1994) for data compression. It was later adapted to the natural language processing (NLP) domain (Sennrich et al., 2016) and has been the favored approach to tokenization for the most popular language models (Grattafiori et al., 2024; Brown et al., 2020) due to its efficiency and robustness to rare words. Byte pair encoding has seen success in representing modalities outside of language, including molecular graphs (Shen and Póczos, 2024), electroencephalogram (EEG) (Klymenko et al., 2023), and more generally, physiological signals.(Tavabi and Lerman, 2021). However, in theses cases the byte pair encoded representations are simply used for classification. Most recently, Tahery et al. (2024) leveraged quantization and BPE to compress ECG signals and pass as inputs to a BERT (Devlin et al., 2019) model for SSL. However, in their work, they only use this representation for classification. As previously mentioned, we believe classification alone may limit aspects of ECG interpretation, thus we utilize these representations for generative diagnosis. Additionally, previous works utilized a pre-existing BPE tokenizer based on SentencePiece (Kudo and Richardson, 2018), and do not conduct further analysis of how the BPE algorithm is merging the ECGs.
3 Methods
This section provides detailed information on the datasets, preprocessing, ECG signal encoding with ECG-Byte, and LLM training for NLG.
3.1 Dataset and Preprocessing
Dataset
In this study, we use variants of the MIMIC-IV ECG (Gow et al., 2023) and PTB-XL datasets (Wagner et al., 2020) for NLG. We use MIMIC-IV ECG pretraining curated by Zhao et al. (2024) that contains question prompts generated by GPT-4o alongside the ECG and clinical notes. Additionally, we use the ECG-QA dataset (Oh et al., 2023), a dataset that uses the ChatGPT API to generate naturalistic, clinically relevant question and answer pairs about the ECG signals from the MIMIC-IV ECG and PTB-XL datasets. The baselines we compare our results with all utilize the single-verify, single-choose, and single-query categorized questions from the ECG-QA dataset. single-verify corresponds to yes or no questions, single-choose corresponds to where a selected answer is made from two given options, and single-query consists of open-ended questions. The ECG signals collected from both both datasets (i.e., MIMIC-IV ECG and PTB-XL) are sampled at 500 Hz for 10 seconds, resulting in a 5000 length, 12 lead ECG.
Preprocessing
We preprocess all datasets used in the study in the same manner to maintain consistency. We first convert the ordering of the leads for the MIMIC-IV ECG dataset (i.e., [‘I’, ‘II’, ‘III’, ‘aVR’, ‘aVF’, ‘aVL’, ‘V1’, ‘V2’, ‘V3’, ‘V4’, ‘V5’, ‘V6’])to the PTB-XL dataset format (i.e., [‘I’, ‘II’, ‘III’, ‘aVL’, ‘aVR’, ‘aVF’, ‘V1’, ‘V2’, ‘V3’, ‘V4’, ‘V5’, ‘V6’]). We then use a notch filter at 50 Hz and 60 Hz to remove powerline interference. Each frequency is targeted with a quality factor of 30 to minimize distortion, and filtering is applied bidirectionally to prevent phase shifts. Next, a fourth-order Butterworth bandpass filter with a range of 0.5–100 Hz isolates relevant ECG components while attenuating high-frequency noise and low-frequency drift. To address baseline wander caused by respiratory or movement artifacts, we bidirectionally apply a fourth-order Butterworth highpass filter with a cutoff of 0.05 Hz. After filtering, we apply wavelet denoising to further reduce noise. Using the Daubechies-6 (db6) wavelet at level 4, we decompose each ECG signal into wavelet coefficients. A soft threshold, based on the median absolute deviation of the detail coefficients, is applied to each coefficient level to suppress noise, ensuring values near zero are excluded from reconstruction. Since 250 Hz is a generally accepted sampling frequency adequate for heartbeat analysis (Kwon et al., 2018), we downsample the 500 Hz sampling frequency to 250. We then segment the 10 second signal to non-overlapping windows of 2 seconds, and use each 12 lead 2 second segment of the ECG signal as input to the model. However, for training the tokenizer, we did not want to introduce this discontinuity across the full 10 seconds. Thus, we utilize the unsegmented, 10 second ECG signal for training ECG-Byte. Lastly, during the unsegmented preprocessing pipeline, we record the global 1st and 99th percentiles out of 300,000 samples to utilize in our later steps of training ECG-Byte for normalization.
3.2 ECG as Bytes
Sampling
Following established practices in NLP (Dagan et al., 2024), we train ECG-Byte on a representative subset of the total dataset, selected using stratified sampling based on morphological clustering. To extract features from each unsegmented ECG, we compute statistical measures, frequency and time domain features, morphological characteristics, and wavelet coefficients. Principal Component Analysis (PCA) (Wold et al., 1987) is applied for dimensionality reduction, retaining 95% of the variance, followed by feature scaling. The optimal number of clusters is determined using the Elbow Method and Silhouette Analysis (Rousseeuw, 1987), with the smaller result chosen. K-means clustering (MacQueen, 1967) is then applied to the scaled PCA-transformed features. If K-means fails to yield distinct clusters, DBSCAN (Ester et al., 1996) is used as a fallback. Stratified sampling is performed by randomly selecting ECGs from each cluster in proportion to its size, resulting in a total sample of 200,000 ECGs.
Quantization
To ensure consistency across ECG signals, we normalize each input by scaling it to a fixed range and encoding it into a symbolic representation. Let $X\in\mathbb{R}^{C\times T}$ denote an ECG signal matrix, where $C$ is the number of ECG leads and $T$ represents the number of sampled time points per lead. In this study, $C=12$ and $T=500$ unless specified otherwise. Let $p_{1}$ and $p_{99}$ represent the 1st and 99th percentiles of $X$ across all leads and time points sampled earlier during preprocessing, respectively. The normalization process is defined as follows:
$X_{\text{norm}}=\frac{X-(p_{1}-\epsilon_{1})}{(p_{99}+\epsilon_{1})-(p_{1}-\epsilon_{1})+\epsilon_{2}}$ (1)
where $\epsilon_{1}=0.5$ is a constant to make up for the sampled percentiles and $\epsilon_{2}=10^{-6}$ is a small constant added to prevent division by zero. This transformation shifts and scales $X$ so that the normalized values fall within the range [0, 1]. We then apply clipping to ensure that values remain strictly within this range:
$X_{\text{clipped}}=\text{clip}(X_{\text{norm}},0,1)$ (2)
Inspired by previous works (Klymenko et al., 2023; Tavabi and Lerman, 2021; Chen et al., 2022; Han et al., 2024), we quantize $X_{\text{clipped}}$ into discrete levels for symbolic representation. Let $\mathcal{A}$ be the set of 26 symbols, corresponding to the lowercased letters in the English alphabet, $\mathcal{A}=\{\text{a},\text{b},\ldots,\text{z}\}$. The alphabet size $|\mathcal{A}|=26$ defines the number of discrete levels. We scale and floor $X_{\text{clipped}}$ to integer values, then take the minimum between the floored value and the maximum number of bins as the following:
$X_{\text{quant}}=\text{min}(\lfloor X_{\text{clipped}}\times|\mathcal{A}|\rfloor,(|\mathcal{A}|-1))$ (3)
Finally, each integer value in $X_{\text{quant}}$ is mapped to a corresponding symbol in $\mathcal{A}$ to yield the symbolic signal, which serves as a discrete representation of the ECG. After transforming each ECG signal instance into its symbolic form, we first flatten each symbolic ECG instance $X_{\text{quant}}^{(i)}$ into a 1-dimensional sequence of symbols:
$X_{\text{symb}}^{(i)}=\text{flatten}(X_{\text{quant}}^{(i)}),\quad X_{\text{symb}}^{(i)}\in\mathcal{A}^{CT}$ (4)
where $i$ indexes over all instances in the dataset, and $X_{\text{symb}}^{(i)}$ is the flattened sequence of symbols of length $C\cdot T$. Next, we concatenate all flattened instances $X_{\text{symb}}^{(1)},X_{\text{symb}}^{(2)},\ldots,X_{\text{symb}}^{(N)}$ across the entire dataset to form a single, long symbolic sequence:
$\mathbf{X}_{\text{concat}}=X_{\text{symb}}^{(1)}\|X_{\text{symb}}^{(2)}\|\cdots\|X_{\text{symb}}^{(N)},\quad\mathbf{X}_{\text{concat}}\in\mathcal{A}^{NCT}$ (5)
[…]
ECG-Byte Training Process
After obtaining the string representation $\mathbf{X}_{\text{concat}}$ of the ECG dataset, we train ECG-Byte to compress the discretized ECG signals by iteratively merging the most frequent byte pairs into single tokens, following the BPE algorithm. The process starts by converting $\mathbf{X}_{\text{concat}}$ into an ID vector of 8-bit unsigned integers and initializing a vocabulary map (vocab) for string representations of bytes and a vocab_tokens map to encode bytes as singleton lists. IDs and vocab are initialized to cover the full byte range (0–255), mapping symbols in $\mathcal{A}$ to ASCII values (97–122), while reserving other byte values for unknown bytes. As merging proceeds, new tokens are assigned unique integer IDs starting from 256, acting as abstract labels for progressively larger token units. For each merge iteration, ECG-Byte calculates adjacent byte pair frequencies using a parallelized get_stats function, efficiently aggregating counts via a fold-and-reduce strategy. The most frequent pair is identified as the ”best pair” to merge, and the merge function replaces occurrences of this pair in the ID vector with a new token ID, extending the vocabulary and updating vocab_tokens accordingly. This process repeats until the specified number of merges is reached or no pairs remain. The output includes the encoded ID vector, the extended vocabulary map, and a history of merge operations. Existing tokenizers, such as SentencePiece (Kudo and Richardson, 2018) or HuggingFace (Wolf et al., 2020), were not used due to their complexity and integration issues, which hindered interpretability. ECG-Byte, implemented in Rust for speed, provides a lightweight, flexible framework for representing ECG signals as discrete tokens while drawing inspiration from HuggingFace’s tokenizer (Wolf et al., 2020). Detailed pseudocode for the main training pipeline is provided in Algorithm 1, with merge and get_stats functions detailed in Appendix A.1.
ECG-Byte Encoding Process
After training ECG-Byte, we encode any quantized ECG signal $X_{\text{symb}}$ by first converting each byte in the ECG signal to a 32-bit unsigned integer and building a trie structure, where each node represents a byte or a merged token sequence from prior encoding steps. The trie is initialized with single-byte tokens (0-255) and is extended with custom token sequences from the learned merge history. For each byte sequence in the input, the encoding function traverses the trie to find the longest match, replacing matched sequences with their assigned token IDs. If no match is found, the byte is added to the output as-is. The final encoded sequence is returned as output_ids, where we will denote as $X_{\text{ID}}$. We provide the detailed pseudocode of the encoding process in Algorithm 2.
Input: Input $X_{concat}$, Number of merges num_merges
Output: Tuple containing final encoded IDs, vocabulary map, and merge history
Function byte_pair_encoding($X_{concat}$, num_merges):
       Convert $X_{concat}$ to ID vector ids by casting each byte to $u32$;
       Initialize vocab with mappings from IDs 00 to $255$ to their string representations;
       Initialize vocab_tokens with mappings from IDs 00 to $255$ to singleton lists;
       Initialize empty list merges;
      for $i\leftarrow 0$ to num_merges exclusive do
             pairs $\leftarrow$ get_stats(ids) using parallel processing;
            if pairs is empty then
                   break
             end if
            best_pair $\leftarrow$ Pair in pairs with highest frequency;
             if best_pair is not found then
                   break
             end if
            new_id $\leftarrow 256+i$;
             ids $\leftarrow$ merge(ids, best_pair, new_id);
            vocab[new_id] $\leftarrow$ concat(vocab[best_pair.0], vocab[best_pair.1]);
            new_token $\leftarrow$ concat(vocab_tokens[best_pair.0], vocab_tokens[best_pair.1]);
             vocab_tokens[new_id] $\leftarrow$ new_token;
            merges.append(new_token, new_id);
       end for
      return (ids, vocab, merges);
Algorithm 1 Training Process for ECG-Byte
Input: Input $X_{\text{symb}}$, Merge history merges containing pairs of token sequences and their token IDs
Output: Vector of encoded IDs
Function encode($X_{\text{symb}}$, merges):
       ids $\leftarrow$ Convert $X_{\text{symb}}$ to vector of u32 by casting each byte;
       Initialize root TrieNode trie_root using TrieNode::new();
      for b $\leftarrow 0$ to $255$ do
             insert(trie_root, [b], b);
       end for
      foreach (token_sequence, token_id) in merges do
             insert(trie_root, token_sequence, token_id);
       end foreach
      Initialize empty list output_ids;
       i $\leftarrow 0$;
      while i $<$ length of ids do
             node $\leftarrow$ trie_root;
             match_len $\leftarrow 0$;
             match_id $\leftarrow$ None;
            for j $\leftarrow$ i to length of ids do
                   id $\leftarrow$ ids[j];
                  if id exists in node.children then
                         node $\leftarrow$ node.children[id];
                        if node.token_id is not None then
                               match_len $\leftarrow$ j - i + 1;
                               match_id $\leftarrow$ node.token_id;
                         end if
                   end if
                  else
                         break
                   end if
             end for
            if match_id is not None then
                   output_ids.append(match_id);
                   i $\leftarrow$ i + match_len;
             end if
            else
                   output_ids.append(ids[i]);
                   i $\leftarrow$ i + 1;
             end if
       end while
      return output_ids;
Algorithm 2 Encoding Process for ECG-Byte
3.3 Large Language Model
In this study, we utilize the Llama-3.2-1B (Grattafiori et al., 2024) checkpoint through the HuggingFace API (Wolf et al., 2020) unless specified otherwise. The Llama 3.2 series model is a variant of the Llama 3 models (Grattafiori et al., 2024) and support context lengths of up to 128k tokens. They are notable for their superior performance despite having 1 billion parameters, making them highly efficient and capable models to test our methodology upon. We also provide an ablation study in subsection 5.4 where we utilize other popular LLMs, such as GPT2 XL 1.5B (Radford et al., 2019), Gemma 2B (Team et al., 2024), and OPT 1.3B (Zhang et al., 2022).
3.4 Learning Objective
The learning objective for training the LLM considers a sequence composed of three parts, $\{X_{\text{ID}},Q,\mathcal{S}\}$, where $X_{\text{ID}}\in\mathcal{V}^{M}$ represents the encoded ECG sequence of length $l_{X_{\text{ID}}}=|X_{\text{ID}}|$, with each token drawn from the extended vocabulary $\mathcal{V}$ of size $M$, $Q$ represents the tokenized question, and $\mathcal{S}$ denotes the tokenized answer sequence. The input sequence includes special tokens: - [BOS] as the beginning-of-sequence token, - [SIG_START] and [SIG_END] to indicate the start and end of the encoded ECG sequence, and - [EOS] as the end-of-sequence token for the generated answer. The motivation for adding [SIG_START] and [SIG_END] special tokens is inspired by Liu et al. (2023), where they utilize special tokens indicating the start and end of the image. Thus, the full input sequence is structured as:
[…]
$L=1+1+l_{X_{\text{ID}}}+1+l_{Q}+l_{\mathcal{S}}+1,$
[…]
$\mathcal{L}_{\text{NLL}}=-\sum_{l^{`}=l_{X_{\text{ID}}}+l_{Q}+4}^{L}\log P(s_{l^{`}}\mid\texttt{Context},s_{<{l^{`}}};\theta),$ (6)
[…]
4 Experiments
4.1 Experimental Settings
We fine-tuned the LLM using the AdamW optimizer (Kingma and Ba, 2017) with a learning rate of $41e-$ 4, weight decay of $21e-$ 2, and a custom learning rate scheduler. This scheduler applies an initial learning rate $\mathrm{init\_lr}$ scaled by the model’s hidden dimension ($\mathrm{d_{model}}^{-0.5}$) and dynamically adjusts it based on training steps, with a warm-up phase of 500 steps. The learning rate at step $\mathrm{n_{steps}}$ is updated as $\text{lr}=\mathrm{init\_lr}\times\min\left(\mathrm{n_{steps}}^{-0.5},\ \mathrm{n_{warmup}}^{-1.5}\times\mathrm{n_{steps}}\right)$. We set $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1e-8$, batch size 2, and trained for 1 epoch. Additionally, we only train on a randomly sampled subset of 400,000 ECG instances for each respective dataset due to computational resources, unless specified otherwise. We also utilize LoRA (Hu et al., 2021) to finetune the LLM with rank = 16, $\alpha_{LoRA}=32$, and dropout = 0.05. We conduct our experiments on 4 NVIDIA RTX A6000 48 GB GPUS.
During inference, we evaluate our model with number of merges num_merges = 3500, sequence length $L=1024$, and ECG length $T=500$ unless specified otherwise. We use popular metrics for NLG namely the BLEU-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Meteor (Banerjee and Lavie, 2005), and BertScore F1 (Zhang et al., 2020) metrics.
[…]
tab:main_results Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1 ECG-Chat (Zhao et al., 2024) MIMIC-IV ECG Pretrain PTB-XL 11.19 29.93 35.10 - $L_{CL}$ 8.10 ± 0.25 31.36 ± 0.31 27.55 ± 0.36 89.35 ± 0.04 $L_{MIM}$ 6.21 ± 0.22 30.63 ± 0.13 24.91 ± 0.14 90.44 ± 0.04 $L_{MERL}$ (Liu et al., 2024) 10.22 ± 0.25 32.95 ± 0.12 25.60 ± 0.17 89.94 ± 0.01 $L_{CL}$ + $L_{MIM}$ 9.33 ± 0.22 30.45 ± 0.21 24.37 ± 0.36 90.29 ± 0.02 ECG-Byte 11.00 ± 0.19 33.41 ± 0.05 24.95 ± 0.09 90.02 ± 0.01 $L_{CL}$ ECG-QA MIMIC-IV ECG-QA MIMIC-IV 10.22 ± 0.06 38.41 ± 0.48 24.66 ± 0.23 90.42 ± 0.09 $L_{MIM}$ 7.90 ± 0.23 29.28 ± 0.38 19.03 ± 0.11 67.91 ± 0.17 $L_{MERL}$ (Liu et al., 2024) 10.95 ± 0.24 38.18 ± 0.58 26.24 ± 0.36 90.80 ± 0.06 $L_{CL}+L_{MIM}$ 8.57 ± 0.14 34.00 ± 0.25 25.22 ± 0.30 87.72 ± 0.04 ECG-Byte 11.23 ± 0.12 42.49 ± 0.53 27.08 ± 0.15 91.30 ± 0.04 $L_{CL}$ ECG-QA PTB-XL ECG-QA PTB-XL 8.89 ± 0.25 28.63 ± 0.47 18.45 ± 0.31 72.63 ± 0.40 $L_{MIM}$ 15.14 ± 0.28 46.71 ± 0.41 29.64 ± 0.30 92.12 ± 0.10 $L_{MERL}$ (Liu et al., 2024) 13.84 ± 0.19 40.14 ± 0.39 26.24 ± 0.35 91.88 ± 0.09 $L_{CL}+L_{MIM}$ 14.72 ± 0.27 42.88 ± 0.13 28.25 ± 0.27 89.40 ± 0.01 ECG-Byte 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
Table 1: NLG mean results with standard deviations over 5 random seeds comparing against different baselines.
5 Results
5.1 Natural Language Generation
We present our main results in Table LABEL:tab:main_results, comparing ECG-Byte with prior works and self-implemented two-stage pretraining methods. Notably, Zhao et al. (2024) is not directly comparable due to differing data splits and pretraining datasets, though reported metrics use the same datasets. Zhao et al. (2024) train on the full MIMIC-IV ECG Pretrain dataset, finetune on an instruction-tuning dataset for ECG-related conversations, and evaluate on PTB-XL (Wagner et al., 2020) using a unified question: “Could you please help me explain my ECG?” To establish baselines, we implement generic two-stage pretraining methods: $L_{CL}$, $L_{MIM}$, and $L_{CL}+L_{MIM}$. Here, $L_{CL}$ employs contrastive learning (Liu et al., 2024; Gopal et al., 2021; Pham et al., 2024; Kiyasseh et al., 2021), $L_{MIM}$ uses Masked Image Modeling (MIM) (Choi et al., 2023; Na et al., 2024; Yang et al., 2022), and $L_{CL}+L_{MIM}$ combines both (Oh et al., 2022; McKeen et al., 2024). These approaches utilize pretrained CLIP (Radford et al., 2021) and ViT (Dosovitskiy et al., 2021), where ECG signals are transformed into three-channel images for finetuning. We adapt Liu et al. (2024)’s state-of-the-art contrastive method ($L_{MERL}$) for fair comparison. Their most effective model uses a 1D ResNet backbone (He et al., 2015); hence, we employ ResNet101 for direct ECG signal processing. For $L_{CL}$, $L_{MIM}$, $L_{CL}+L_{MIM}$, and $L_{MERL}$, training is conducted on the full, preprocessed MIMIC-IV ECG dataset with a batch size of 64 during the first stage. Implementation details for both training stages are in Appendix B. Table LABEL:tab:main_results demonstrates ECG-Byte’s effectiveness, showing competitive or superior performance across all metrics and datasets compared to other methods. Qualitative examples are provided in Appendix C.2.
5.2 Cross Dataset Transferability
We present the results of cross-dataset transferability in Table LABEL:tab:cross, comparing our approach, ECG-Byte, with two-stage pretraining methods. ECG-Byte achieves the best zero-shot transfer performance from the ECG-QA PTB-XL dataset to the ECG-QA MIMIC-IV dataset. When transferring from the ECG-QA MIMIC-IV dataset to the ECG-QA PTB-XL dataset, although other 2-stage pretraining methods demonstrate higher performance, ECG-Byte maintains competitive results across all metrics.
5.3 Efficiency of ECG-Byte
We compare the efficiency of our end-to-end approach using ECG-Byte with 2-stage pretraining methods in Table LABEL:tab:efficiency. First, we examine the amount of data required for each method. As previously noted, the first stage of the two-stage pretraining methods is trained on the full MIMIC-IV ECG dataset (Johnson et al., 2023) using segmented ECGs, which are also used as input during the second stage. While ECG-Byte is trained on unsegmented ECGs, we convert the number of unsegmented ECGs to an equivalent count of segmented ECGs. Additionally, the reduced data requirement for ECG-Byte is due to our sampling approach, where only a subset of ECGs is used to train the tokenizer. Under these settings, our proposed method achieves competitive results using approximately $\sim$48of the data required for 2-stage pretraining methods. In terms of training time, our approach requires less than half the time needed for two-stage pretraining. The training time for the two-stage methods is averaged across our self-implemented approaches ($L_{CL}$, $L_{MIM}$, and $L_{CL}+L_{MIM}$) and the $L_{MERL}$ method proposed by Liu et al. (2024).
[…]
tab:cross Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1 $L_{CL}$ ECG-QA MIMIC-IV ECG-QA PTB-XL 11.64 ± 0.45 41.48 ± 0.11 25.74 ± 0.13 91.24 ± 0.05 $L_{MIM}$ 11.70 ± 0.29 42.22 ± 0.28 26.41 ± 0.10 91.51 ± 0.03 $L_{MERL}$ (Liu et al., 2024) 11.53 ± 0.19 39.23 ± 0.40 25.58 ± 0.28 91.59 ± 0.03 $L_{CL}$ + $L_{MIM}$ 9.71 ± 0.10 35.10 ± 0.28 24.91 ± 0.19 87.88 ± 0.08 ECG-Byte 8.70 ± 0.04 40.39 ± 0.40 23.29 ± 0.18 91.51 ± 0.03 $L_{CL}$ ECG-QA PTB-XL ECG-QA MIMIC-IV 5.10 ± 0.04 22.77 ± 0.28 14.63 ± 0.32 77.89 ± 0.13 $L_{MIM}$ 7.68 ± 0.46 35.77 ± 0.13 22.32 ± 0.33 90.28 ± 0.07 $L_{MERL}$ (Liu et al., 2024) 7.39 ± 0.15 28.33 ± 0.58 18.59 ± 0.35 89.30 ± 0.05 $L_{CL}$ + $L_{MIM}$ 7.49 ± 0.21 30.53 ± 0.59 20.25 ± 0.27 86.53 ± 0.11 ECG-Byte 7.86 ± 0.13 35.01 ± 0.41 21.49 ± 0.24 90.29 ± 0.07
Table 2: Mean results with standard deviations over 5 random seeds on zero shot cross-dataset transferability.
[…]
tab:efficiency 1st Stage 2nd Stage ECG-Byte end-to-end # of Data 2,513,435 400,000 1,000,000 400,000 Total # of Data 2,913,435 1,400,000 Time (minutes) $\sim$1258.50 $\sim$469.25 $\sim$385.12 $\sim$420.32 Total Time (minutes) $\sim$1727.75 $\sim$805.44
Table 3: Efficiency of our method compared against 2-stage pretraining methods.
5.4 Ablation Study
We conduct several ablation studies to show the variability of performance with ECG-Byte when we alter the LLM used for finetuning, use different sequence lengths $L$ when inputting to the LLM, training ECG-Byte with various number of merges num_merges, and varying ECG lengths $T$. With the exception of the ablating parameter, we fix all other parameters to num_merges = 3500, $L=1024$, and $T=500$. We report results on the test set of the PTB-XL variant of ECG-QA (Oh et al., 2023) unless specified otherwise.
[…]
tab:diff_llm LLM BLEU-4 Rouge-L Meteor BertScore F1 GPT2 XL 1.5B (Radford et al., 2019) 12.30 ± 0.19 41.33 ± 0.57 26.48 ± 0.33 92.00 ± 0.06 Gemma 2B (Team et al., 2024) 13.78 ± 0.18 45.48 ± 0.55 28.32 ± 0.23 92.01 ± 0.02 OPT 1.3B (Zhang et al., 2022) 12.26 ± 0.20 41.84 ± 0.52 26.21 ± 0.29 91.78 ± 0.04 Llama 3.2 1B (Grattafiori et al., 2024) 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
Table 4: Ablation study on using different LLMs.
Different LLMs
We show the variability in performance of ECG-Byte when using different LLMs with similar numbers of parameters in Table LABEL:tab:diff_llm. While Llama 3.2 1B (Grattafiori et al., 2024) achieves the best results, GPT2 XL 1.5B (Radford et al., 2019), Gemma 2B (Team et al., 2024), and OPT 1.3B (Zhang et al., 2022) also deliver comparable performances. These findings demonstrate that our method is not limited to Llama 3.2 1B but can achieve similar results across a variety of LLMs.
Sequence Length
Input lengths for LLMs are an important parameter to consider for efficient training since the calculation of attention is quadratic with respect to the input length (Vaswani et al., 2023). We present results on different sequence lengths $L$ in Table LABEL:tab:seq_len. Although the difference is not substantial, we can see that when $L=1024$ and $L=2048$ the model yields higher performance then $L=512$. We attribute this to the rate of truncation and padding for the encoded ECG. Observing Figure 2, most ECGs were being encoded to token sequence lengths of around 500 to 1500. Therefore, we hypothesize that when $L=512$ a large portion of the ECG token sequence gets truncated, resulting in lower performance.
[…]
tab:seq_len $L$ BLEU-4 Rouge-L Meteor BertScore F1 512 13.61 ± 0.15 48.15 ± 0.57 29.10 ± 0.28 92.41 ± 0.05 1024 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07 2048 13.88 ± 0.22 45.21 ± 0.48 28.31 ± 0.27 90.88 ± 0.02
Table 5: Ablation study on varying sequence lengths $L$.
Number of Merges
The number of merges num_merges performed during training ECG-Byte corresponds to how much the algorithm compresses the concatenated sequence of quantized ECGS $\mathbf{X}_{\text{concat}}$. More num_merges means more compression, which can affect the expressiveness of the encoded sequence. In Table LABEL:tab:num_merges, we show the performance of our method with different num_merges. The results indicate that while performance varies slightly with the number of merges, values of num_merges greater than 500 generally yield similar outcomes.
[…]
tab:num_merges num_merges BLEU-4 Rouge-L Meteor BertScore F1 500 13.61 ± 0.53 46.50 ± 0.28 28.49 ± 0.49 92.33 ± 0.02 1750 14.50 ± 0.25 46.74 ± 0.48 30.03 ± 0.25 92.55 ± 0.01 2500 15.10 ± 0.39 46.37 ± 0.28 30.12 ± 0.23 92.53 ± 0.05 3500 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
Table 6: Ablation study on varying number of merges num_merges.
ECG length
Lastly, we show the effect of the length $T$ being considered when encoding the ECG with ECG-Byte in Table LABEL:tab:len_ecg. We want to note that for the results of $T=2000$, the full unsegmented ECG is utilized. Consequently, the number of instances available is less than the targeted dataset size of 400,000 (i.e., 97,244). Thus, when $T=2000$, we use the full dataset to train the model. For shorter segment lengths, such as $T=250$ and $T=500$, the model demonstrates strong performances indicating that shorter segments can effectively preserve relevant information for NLG. Interestingly, for $T=2000$, the model achieves the highest performance across all metrics. This suggests that when the model is trained with the full 10 second encoded ECG, it benefits from richer contextual information present in the complete ECG waveform.
[…]
tab:len_ecg $T$ BLEU-4 Rouge-L Meteor BertScore F1 250 12.64 ± 0.20 47.31 ± 0.26 27.97 ± 0.21 92.32 ± 0.06 500 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07 1250 11.01 ± 0.19 43.84 ± 0.28 25.49 ± 0.20 93.07 ± 0.03 2000 14.54 ± 0.17 48.03 ± 0.27 32.11 ± 0.22 92.91 ± 0.04
Table 7: Ablation study on varying lengths $T$.
5.5 ECG-Byte Analysis
[…]
Token Usage and Length Distribution
[…]
We examine the token usage and length distributions for ECG-Byte with num_merges = 3500 on a subsample of 277,840 ECGs from the PTB-XL dataset. The left panel of Figure 2 displays the token usage distribution, showing token frequency (y-axis) ranked in descending order (x-axis). A small subset of tokens dominates the occurrences, while the rest are infrequently used—a typical characteristic of BPE-based tokenization, where common patterns are compressed into frequent tokens and rare patterns into infrequent ones. The right panel of Figure 2 illustrates the token length distribution of the encoded ECGs, with most falling between 500 and 1000 tokens, demonstrating ECG-Byte’s effective compression of the original signal. Additional examples of these distributions are provided in Appendix A.3.
Token to ECG Mapping
Figure 3: A mapping between tokens used for a given ECG Lead I. More examples are provided in Appendix A.2.
To illustrate how ECG-Byte encodes ECG signals, we analyze the mapping between tokens and signal features. Figure 3 shows an example Lead I ECG signal with unique token IDs (represented by different colors) overlayed. The P wave, QRS complex, and T wave are distinctly captured by different tokens, though this precision varies across instances. As demonstrated, ECG-Byte effectively merges key regions of the signal. Additional examples are provided in Appendix A.2 due to page limitations.
Attention Visualizations
Figure 4 visualizes attention weights across a selected ECG lead and text portions of the input after training. We focus on one lead due to the uniformity of attention across encoded signal tokens. For interpretability, the reversed ECG signal is overlayed on the encoded ECG. The model primarily attends to the textual portion of the input sequence, as shown in Figure 4. Previous studies have debated whether attention visualizations are inherently explainable (Jain and Wallace, 2019; Wiegreffe and Pinter, 2019) and explored their role in vision-language models (Aflalo et al., 2022; Woo et al., 2024; Arif et al., 2024; Cui et al., 2024). These works often observe minimal attention to visual input, with models relying primarily on text. We hypothesize that a similar phenomenon occurs in Figure 4, as the ECG tokens, though represented like text, are 1) newly introduced and 2) perceived as a different modality (e.g., vision). Additional examples are provided in Appendix A.4.
6 Discussion and Conclusion
In this study, we introduce ECG-Byte, a custom BPE algorithm to encode ECGs into a discrete sequence of tokens for conditional autoregressive NLG. ECG-Byte introduces a paradigm shift in generative ECG language modeling by enabling efficient end-to-end training, compared to traditional two-stage pretraining approaches. Our pipeline demonstrates strong performances, achieving results comparable to two-stage methods while requiring only half the training time and approximately $48\%$ of the data. In addition to its efficiency, ECG-Byte enhances interpretability.
Figure 4: The attention weight overlayed on top of both the text (top) and ECG (bottom). More examples are provided in Appendix A.4.
By analyzing its underlying mechanism, we observe that critical ECG regions, such as the P wave, the QRS complex, and the T wave, are effectively grouped during tokenization, as illustrated in Figure 3. Furthermore, the reversibility of the compressed token sequence allows us to trace each token back to its original ECG signal segment, providing insight into the specific portions of the signal attended to by the model. However, as shown in Figure 4, the model’s attention weight distribution resembles that of vision language models, focusing primarily on the textual components of the input sequence during generation.
This work is in its early stages and needs further exploration. Future directions include: (1) refining BPE merging rules to better capture ECG-specific features, (2) adopting more advanced quantization techniques that preserve time-series characteristics (Carson et al., 2024b; Elsworth and Güttel, 2020; Carson et al., 2024a), (3) introducing stronger modality-specific distinctions, such as embeddings beyond [SIG_START] and [SIG_END] (Gui et al., 2023), and (4) extending ECG-Byte for conversational tasks through instruction tuning.
Limitations
One of the main limitations of this work is the scale in terms of computing and data. Since we only used a subset of the data to train and test our method, we were unable to train the model to its full potential. However, even with only using a small subset of the data, we are able to see extremely promising results compared to other 2-stage SSL pretraining methods. Therefore, we do not view this limitation as a major bottleneck.
[…]
This work is done in collaboration with the Mario Lemieux Center for Heart Rhythm Care at Allegheny General Hospital. We thank Wenhao Ding, Haohong Lin, Shiqi Liu, and Hyoeun Kang for the valuable discussions.
Appendix A ECG-Byte
A.1 Additional Pseudocode for ECG-Byte
Input: Array of IDs ids, Pair to merge pair as $(u_{1},u_{2})$, New ID new_id
Output: Merged vector of IDs
Function merge(ids, pair, new_id):
       Initialize empty vector new_ids with capacity of ids;
       i $\leftarrow$ 0;
       while i $<$ length of ids do
[…]
                   new_ids.append(new_id);
                   i $\leftarrow\texttt{i}+2$;
             end if
            else
                   new_ids.append(ids[i]);
                   i $\leftarrow\texttt{i}+1$;
             end if
       end while
      return new_ids;
Algorithm 3 Merging a pair in an ID array
Input: Array of IDs ids
Output: HashMap of pairs and their frequencies
Function get_stats(ids):
       pair_counts $\leftarrow$ Parallel fold operation:;
       foreach window of size 2 in ids do
             Let $(u_{1},u_{2})\leftarrow$ elements of the window;
             Increment the count of $(u_{1},u_{2})$ in local pair_count;
       end foreach
      pair_counts $\leftarrow$ Parallel reduce operation to combine local pair_count HashMaps;
       return pair_counts;
Algorithm 4 Calculating Frequency of Byte Pairs in an Array
A.2 Mapping between Token and ECG
We add more examples of the mapping between the ECG signal and the encoded tokens for ECG-Byte in Figures 5 and 6.
Figure 5: A mapping between tokens used for a given ECG Leads I, II, III, aVL, aVR, aVF.
Figure 6: A mapping between tokens used for a given ECG Leads V1, V2, V3, V4, V5, V6.
A.3 Token usage and length distribution for varying num_merges
We add more examples of the token usage and length distributions for varying num_merges in Figure 7.
Figure 7: Plots of the token usage and length distributions for ECG-Byte where num_merges is 500, 1750, and 2500 from top to bottom.
A.4 Attention Visualizations
We add more visualizations of the attention weights in Figures 8, 9, 10, 11, 12, 13, 14, 15.
Figure 8: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 9: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 10: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 11: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 12: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 13: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 14: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 15: The attention weight overlayed on both text (top) and ECG (bottom).
Appendix B 2-stage Pretraining Approaches
To be consistent, we normalize each ECG in the same manner as described in subsection 3.2. Consider a dataset of $N$ ECG-image and clinical note pairs, denoted as $\{(I_{i},O_{i})\}_{i=1}^{N}$, where: $I_{i}\in\mathbb{R}^{3\times C\times T}$ is the $i$-th normalized and replicated ECG image, obtained by stacking the clipped ECG signal $X_{\text{clipped}}$ along the channel dimension: $I_{i}=\text{stack}(X_{\text{clipped}},X_{\text{clipped}},X_{\text{clipped}})$. The reason we do this is because we need to create RGB images to use pretrained image models like ViT (Dosovitskiy et al., 2021) and CLIP (Radford et al., 2021).
$O_{i}$ is the corresponding clinical note for the $i$-th ECG, serving as the textual description. Note that $O_{i}$ differs from $S$ in the autoregressive setup, where $S$ represents the tokenized answer sequence provided by either ECG-QA (Oh et al., 2023) or MIMIC-IV ECG pretraining (Zhao et al., 2024).
Given these two features $I$ and $O$ we then describe the contrastive, masked, and dual approaches implemented for our baselines that are derived from commonly used techniques used throughout previous works (Oh et al., 2022; Choi et al., 2023; McKeen et al., 2024; Pham et al., 2024; Tang et al., 2024a, b; Vaid et al., 2022).
B.1 Contrastive learning approaches
We utilize a pretrained CLIP Radford et al. (2021) checkpoint, namely ‘openai/clip-vit-base-patch32’, provided by HuggingFace (Wolf et al., 2020) to encode ECG signals $I$ and text labels $O$ into a shared embedding space. Let $f_{\text{img}}:\mathbb{R}^{3\times C\times T}\to\mathbb{R}^{d}$ and $f_{\text{txt}}:\text{Text}\to\mathbb{R}^{d}$ be the image and text encoders of the pretrained CLIP model, respectively. The embeddings for the $i$-th pair are computed as:
$z_{i}^{\text{img}}=f_{\text{img}}(I_{i}),\quad z_{i}^{\text{txt}}=f_{\text{txt}}(O_{i}),$
where $z_{i}^{\text{img}},z_{i}^{\text{txt}}\in\mathbb{R}^{d}$. The CLIP loss function $\mathcal{L}_{\text{CLIP}}$ aligns the embeddings of corresponding ECG signals and text labels while contrasting them with non-matching pairs. This is formulated as:
$\mathcal{L}_{\text{CL}}=-\frac{1}{N}\sum_{i=1}^{N}\Bigg{[}$ $\log\frac{\exp\left(\text{sim}(z_{i}^{\text{img}},z_{i}^{\text{txt}})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(z_{i}^{\text{img}},z_{j}^{\text{txt}})/\tau\right)}$
$+$ $\log\frac{\exp\left(\text{sim}(z_{i}^{\text{txt}},z_{i}^{\text{img}})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(z_{i}^{\text{txt}},z_{j}^{\text{img}})/\tau\right)}\Bigg{]}$
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, and $\tau$ is a learnable temperature parameter.
To integrate the pretrained CLIP model into our language model for joint reasoning over ECG signals and text, we project the frozen image embeddings $z_{i}^{\text{img}}$ into the language model’s hidden space. Let $W\in\mathbb{R}^{h\times d}$ be a learnable projection matrix, where $h$ is the hidden dimension of the language model. The projected embeddings are:
$z_{i}^{\text{clip}}=Wz_{i}^{\text{img}}.$
[…]
B.2 Masked image modeling approaches
Consider the normalized ECG image $I\in\mathbb{R}^{3\times C\times T}$ obtained as previously described. We utilize a pretrained Vision Transformer (ViT) model (Dosovitskiy et al., 2021), specifically the ‘google/vit-base-patch16-224-in21k’ checkpoint provided by HuggingFace (Wolf et al., 2020).
The image $I$ is partitioned into $P$ non-overlapping patches. Let $N$ be the number of images in our dataset, and $I_{i}$ denote the $i$-th image. The ViT encoder $f_{\text{vit}}$ projects these patches into latent embeddings:
$z_{i}^{\text{patch}}=f_{\text{vit}}(I_{i})\in\mathbb{R}^{P\times d},$
where $d$ is the embedding dimension of the ViT model.
During training, we randomly mask a subset of patches for each image $I_{i}$, creating a binary mask $M_{i}\in\{0,1\}^{P}$, where $M_{i,j}=1$ if patch $j$ is masked and $M_{i,j}=0$ otherwise. The masked embeddings $z_{i}^{\text{masked}}$ are formed by replacing the embeddings of masked patches with a mask token. A reconstruction head $f_{\text{rec}}$ is then applied to predict the pixel-level content of the masked patches:
$\hat{I}_{i}=f_{\text{rec}}(z_{i}^{\text{masked}})\in\mathbb{R}^{P\times d}.$
The masked image modeling loss $\mathcal{L}_{\text{MIM}}$ is computed as the mean squared error (MSE) between the reconstructed embeddings $\hat{I}_{i}$ and the original embeddings $z_{i}^{\text{patch}}$ at the masked positions:
$\mathcal{L}_{\text{MIM}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\sum_{j=1}^{P}M_{i,j}}\sum_{j=1}^{P}M_{i,j}\left\|\hat{I}_{i}[j]-z_{i}^{\text{patch}}[j]\right\|_{2}^{2}.$ (7)
To integrate the MIM representations into the language model for joint reasoning over ECG signals and textual questions, we project the frozen ViT embeddings $z_{i}^{\text{img}}\in\mathbb{R}^{d}$ into the language model’s hidden space. Let $W\in\mathbb{R}^{h\times d}$ be a learnable projection matrix, where $h$ is the hidden dimension of the language model. The projected embeddings are given by:
$z_{i}^{\text{vit}}=Wz_{i}^{\text{img}}.$
[…]
B.3 Dual approaches
The dual approach follows the previous two contrastive and masked image modeling approaches for pretraining the ECG encoder but simply just combines the losses like so:
$\mathcal{L}_{\text{Dual}}=\lambda_{1}\mathcal{L}_{\text{MIM}}+\lambda_{2}\mathcal{L}_{\text{CL}}$
where $\lambda_{1}=\lambda_{2}=1$ in our study.
However, when training the autoregressive LLM, we project both embeddings, $z_{i}^{\text{vit}}$ and $z_{i}^{\text{clip}}$, outputted by their respective frozen encoders via a learnable projection matrix into the language model’s hidden space of dimension $h$. We then concatenate the projected embeddings and pass them through a fusion network to obtain the fused visual embedding $z_{i}^{\text{fused}}\in\mathbb{R}^{h}$:
$z_{i}^{\text{fused}}=f_{\text{fusion}}(\text{concat}(z_{i}^{\text{vit}};z_{i}^{\text{clip}})),$
[…]
Appendix C Additional Results
C.1 Does Larger LLMs Yield Higher Performance?
We present the results of ablating the size of the LLM in Table LABEL:tab:large_llm. Interestingly, the performance across the three different model sizes (1B, 3B, 8B) remains fairly similar. We believe that the limited dataset size prevents the larger models from realizing their full performance potential. We hypothesize that increasing the amount of training data would enable the larger models to leverage their greater capacity, resulting in observable performance improvements.
[…]
tab:large_llm LLM BLEU-4 Rouge-L Meteor BertScore F1 Llama 3.2 1B (Grattafiori et al., 2024) 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07 Llama 3.2 3B (Grattafiori et al., 2024) 14.80 ± 0.17 46.55 ± 0.21 29.53 ± 0.16 92.42 ± 0.01 Llama 3.1 8B (Grattafiori et al., 2024) 13.80 ± 0.16 46.29 ± 0.25 28.56 ± 0.11 92.44 ± 0.05
Table 8: Ablation study on how larger LLMs perform for NLG.
C.2 Qualitative NLG Examples
We provide qualitative NLG examples of successful (Figure 17) and unsuccessful generations (Figure 16).
Figure 16: Randomly sampled NLG results of unsuccessful generations on the PTB-XL test set from ECG-QA.
Figure 17: Randomly sampled NLG results of successful generations on the PTB-XL test set from ECG-QA.

### v2

@@
 Abstract
-Large Language Models (LLMs) have shown remarkable adaptability across domains beyond text, specifically electrocardiograms (ECGs).
-More specifically, there is a growing body of work exploring the task of generating text from a multi-channeled ECG and corresponding textual prompt.
-Current approaches typically involve pretraining an ECG-specific encoder with a self-supervised learning (SSL) objective and using the features output by the pretrained encoder to finetune a LLM for natural language generation (NLG).
-However, these methods are limited by 1) inefficiency from two-stage training and 2) interpretability challenges with encoder-generated features.
-To address these limitations, we introduce ECG-Byte, an adapted byte pair encoding (BPE) tokenizer pipeline for autoregressive language modeling of ECGs.
-This approach compresses and encodes ECG signals into tokens, enabling end-to-end LLM training by combining ECG and text tokens directly, while being much more interpretable since the ECG tokens can be directly mapped back to the original signal.
-Using ECG-Byte, we achieve competitive performance in NLG tasks in only half the time and $\sim$48of the data required by two-stage approaches.
-Data and Code Availability
-This paper uses the ECG-Chat pretraining (Zhao et al., 2024) and ECG-QA datasets (Oh et al., 2023), which were both created from the MIMIC-IV ECG (Johnson et al., 2023) and PTB-XL datasets (Wagner et al., 2020).
-More details about the datasets are provided in Section 3.
-The ECG-Chat pretraining, ECG-QA, MIMIC-IV ECG, and PTB-XL datasets are all freely available on https://github.com/YubaoZhao/ECG-Chat, https://github.com/Jwoo5/ecg-qa, and https://physionet.org/ respectively.
-Lastly, we have released the code at the following link: https://github.com/willxxy/ECG-Byte.
-Institutional Review Board (IRB)
-All datasets used in this study are directly taken from the publicly available, de-identified MIMIC-IV ECG (Johnson et al., 2023) and PTB-XL (Wagner et al., 2020) datasets, thus not requiring IRB approval.
+Large Language Models (LLMs) have demonstrated exceptional versatility across domains, including applications to electrocardiograms (ECGs).
+A growing body of work focuses on generating text from multi-channeled ECG signals and corresponding textual prompts.
+Existing approaches often involve a two-stage process: pretraining an ECG-specific encoder with a self-supervised learning (SSL) objective, followed by finetuning an LLM for natural language generation (NLG) using encoder-derived features.
+However, these methods face two key limitations: inefficiency due to multi-stage training and challenges in interpreting encoder-generated features.
+To overcome these issues, we propose ECG-Byte, an adapted byte pair encoding (BPE) tokenizer pipeline for autoregressive language modeling of ECGs.
+ECG-Byte compresses and encodes ECG signals into tokens, enabling direct end-to-end LLM training by combining ECG and text tokens.
+This approach enhances interpretability, as ECG tokens can be directly mapped back to the original signals.
+Leveraging ECG-Byte, we achieve competitive NLG performance while training 3 times faster and using just 48% of the data required by traditional two-stage methods.
+All code is available at github.com/willxxy/ECG-Byte.
 1 Introduction
@@
 However, accurate ECG analysis is limited in places where ECG expertise is not accessible, exacerbated by the decline and lack of available cardiac electrophysiologists especially in rural areas (Johnson, 2024).
-Figure 1: Comparison of traditional methods and our method on ECG language modeling.
-Traditional methods comprises 2 training stages.
-The first stage aims to learn a good representation of the 12-lead ECG by training an ECG-specific encoder with a self-supervised learning objective via a combinatorial or individual usage of contrastive learning ($L_{CL}$) on the hidden states of the ECG $z_{e}$ and text $z_{t}$ or masked image modeling ($L_{MIM}$).
-The second stage applies the trained encoder to an ECG, which is input alongside a text prompt to a LLM for generation.
-In contrast, our method is end-to-end and can directly train a LLM for generation utilizing ECG-Byte.
+Figure 1: Comparison of traditional and our approach for ECG language modeling.
+Traditional methods follow a two-stage process: (i) training a 12-lead ECG encoder with self-supervised objectives—contrastive learning ($L_{CL}$ between ECG ($z_{e}$) and the textual diagnostic report ($z_{o}$)) and/or masked image modeling ($L_{MIM}$)—to learn robust latent features; and (ii) mapping these ECG features ($z^{\prime}_{e}$) to a shared representation space via a learnable projection layer, then concatenating them with textual prompt vectors ($z_{t}$) as input to a large language model (LLM) for generation.
+In contrast, our method trains an LLM end-to-end using ECG-Byte as an ECG tokenizer.
 The aforementioned facts calls attention to the need for accessible, accurate, and efficient automation of ECG analysis through deep learning.
-Deep learning has reached expert level performance in certain tasks for CVD detection using ECGs (Rajpurkar et al., 2017; Hannun et al., 2019; Qiu et al., 2023b).
-However, most previous works in this domain have succumbed to a crude classification of hard CVD labels (Choi et al., 2023; Nonaka and Seita, 2020; Martin et al., 2021; Strodthoff et al., 2021).
-A problem with this approach is that ECGs often do not exclusively fall into one diagnostic category, instead, there may be many soft labels annotated by expert physicians and the accumulation of these soft labels allow a more detailed, nuanced, and clinically useful interpretation of the ECG.
-The recent onset of Large Language Models (LLMs) provides an opportunity to take a softer, generative, and consequently more flexible approach to ECG analysis.
-There are some recent works that have taken advantage of this approach to ECG analysis (Qiu et al., 2023a; Tang et al., 2024b; Wan et al., 2024; Fu et al., 2024; Zhao et al., 2024).
-A commonality among these works is that they treat multi-channel ECGs similarly to images; they first pretrain an encoder specifically for ECGs with some self-supervised learning (SSL) objective, then apply the learned features to finetune a LLM for natural language generation (NLG).
-However, we observed 2 limitations with this approach: 1) inefficiencies from two-stage training and 2) interpretability challenges with encoder generated features.
-Pretraining a good ECG specific encoder can be a significant computational burden due to large datasets, model size, and long training times.
-Additionally, the latent feature vector output by the ECG encoder cannot be mapped back to the signal, making interpretability difficult when utilizing this feature vector for downstream tasks.
-In this study, we introduce ECG-Byte, an adapted byte pair encoding (BPE) (Gage, 1994) tokenizer designed for end-to-end training for generative ECG language modeling.
-Inspired by prior works demonstrating the effectiveness of creating discrete tokens from continuous values (Chen et al., 2022; Han et al., 2024), we leverage quantization to represent amplitude ranges as discrete symbols.
-Using discrete symbols, we obtain string representations of the ECGs to train ECG-Byte.
-We then apply ECG-Byte to directly finetune a LLM for conditional autoregressive NLG, where the text output is conditioned on the text prompt and tokenized signal.
-We found that our end-to-end training approach is competitive in conditional NLG with only half the time and $\sim$48of the data required by 2-stage pretraining approaches.
-Additionally, since the encoded ECG can be reverse transformed back to the original signal, we can interpret token-level attention-based visualizations, whereas it would be impossible to reverse the hidden latent feature vector outputs by a pretrained encoder.
-Our contributions are summarized as the following:
-  1. 1.
-    We present ECG-Byte, an adapted BPE tokenizer for end-to-end training on autoregressive, conditional NLG.
-  2. 2.
-    We empirically show the efficiency of our method and present competitive performance compared to conventional 2-stage pretraining approaches, proposing a new paradigm for conditional NLG with ECGs.
-  3. 3.
-    We conduct an interpretability study on both ECG-Byte and the LLM by respectively observing how ECG-Byte is merging ECG signals and by utilizing attention visualizations to observe how the LLM is processing ECGs and text.
+Deep learning has reached expert level performance in certain tasks for CVD detection using ECGs (Rajpurkar et al., 2017; Hannun et al., 2019).
+However, most previous works have succumbed to a crude classification of hard CVD labels (Nonaka and Seita, 2020; Martin et al., 2021; Strodthoff et al., 2021).
+A problem with this approach is that ECGs often do not exclusively fall into one diagnostic category, instead, there may be many soft labels annotated by expert physicians and the accumulation of these soft labels allow a more detailed, nuanced, and clinically useful interpretation of the ECG (Singstad and Muten, 2022; Yoo et al., 2021).
+The recent emergence of Large Language Models (LLMs) enables a generative, flexible approach to ECG analysis.
+Recent works have adopted this method (Tang et al., 2024b; Zhao et al., 2024), treating multi-channel ECGs as images by first pretraining an ECG-specific encoder with a self-supervised learning (SSL) objective and then finetuning an LLM for natural language generation (NLG).
+However, this two-stage process has two key drawbacks: (1) training inefficiencies, as pretraining an effective encoder demands significant computational resources due to large datasets, model sizes, and long training times, and (2) interpretability challenges, since the encoder’s latent feature vectors cannot be mapped back to the original signal.
+In this study, we introduce ECG-Byte, a byte pair encoding (BPE) tokenizer (Gage, 1994) designed for end-to-end training of generative ECG language models.
+Building on prior work that converts continuous values to discrete tokens (Chen et al., 2022; Han et al., 2024), we use quantization to represent amplitude ranges as discrete symbols and obtain string representations of ECGs.
+We then directly finetune an LLM for autoregressive natural language generation (NLG), conditioning text output on both the text prompt and tokenized signal.
+Our end-to-end approach is competitive for conditional NLG, with 3 times faster total training time and approximately 48% of the data compared to two-stage pretraining methods.
+Moreover, since the tokenized ECG can be directly traced back to the original signal, token-level attention-based visualizations become interpretable.
+Our contributions are summarized below:
+1. We introduce ECG-Byte, a tokenizer that compresses discretized symbolic representations of ECGs using BPE, enabling direct end-to-end training without the need for an ECG-specific encoder.
+2. We empirically show the efficiency of our method and present competitive performances against conventional two-stage pretraining approaches, proposing a new paradigm for conditional NLG with ECGs.
+3. We perform an interpretability study on ECG-Byte and the LLM, examining how ECG-Byte merges ECG signals and using attention visualizations to analyze how the LLM processes ECGs and text.
+Generalizable Insights about Machine Learning in the Context of Healthcare
+The motivation and generalizable insights of ECG-Byte derive from Vision-Language Models (VLMs) (Liu et al., 2023, 2024b), where image encoders are pretrained on large-scale data and later applied for visual understanding tasks.
+While recent ECG research has focused on developing specialized encoders for classification (Na et al., 2024; Liu et al., 2024a; Kiyasseh et al., 2021) and some works like ECG-Chat (Zhao et al., 2024) have shown potential for using these encoders in NLG, ECG-Byte demonstrates that a simpler approach can achieve competitive results.
+Our key insight is that compressing ECG data via a rule-based method (BPE tokenization) eliminates the need for complex encoder training pipelines that often involve additional overhead due to the size of the encoder, multiple loss objectives and carefully guided feature extraction (Gopal et al., 2021; Oh et al., 2022).
+Our analysis reveals that ECG-Byte effectively preserves clinically relevant signal components (P wave, QRS complex, T wave) within its tokenization scheme.
+Beyond cardiovascular applications, this successful adaptation of BPE for time series data suggests a promising approach for processing other physiological signals such as electroencephalograms (EEGs), photoplethysmograms (PPGs), or electromyograms (EMGs) in generative tasks with LLMs.
 2 Related Works
 2.1 Deep learning for ECGs
-There has been a plethora of works utilizing deep learning for processing ECGs for classification (Rajpurkar et al., 2017; Hannun et al., 2019; Choi et al., 2023; Nonaka and Seita, 2020; Martin et al., 2021; Strodthoff et al., 2021).
+There has been a plethora of works utilizing deep learning for processing ECGs for classification (Rajpurkar et al., 2017; Hannun et al., 2019; Nonaka and Seita, 2020; Martin et al., 2021; Strodthoff et al., 2021).
 Most of these works utilize either convolutional neural networks (CNNs) (Rajpurkar et al., 2017) or transformers (Choi et al., 2023) and exhibit excellent performance at classification tasks.
-There have also been some efforts to frame classification as a retrieval task in order to recover cases similar to the given ECG.(Tang et al., 2024b; Qiu et al., 2023b).
-However, the retrieval approach may struggle with rare or unique ECG patterns that lack good matches in the existing database, potentially leading to missed or inaccurate diagnoses for unusual cases.
+There have also been some efforts to frame classification as a retrieval task in order to recover cases similar to the given ECG (Tang et al., 2024b; Qiu et al., 2023b).
+However, the retrieval approach may struggle with rare or unique ECG patterns that lack good matches in the existing database.
 Additionally, both classification and retrieval tasks may be crude formulations of processing ECGs, since ECGs typically exhibit many characteristics of overlapping CVDs.
@@
 Generative Large Language Models (LLMs) have given the opportunity to take a softer and more clinically similar approach in processing ECGs by generating physician-vetted clinical statements (Qiu et al., 2023a; Tang et al., 2024b; Wan et al., 2024; Fu et al., 2024; Zhao et al., 2024).
-The representation of ECG data has been largely considered in previous works which feed the raw signal into an encoder to obtain a latent representation of the ECG data, which then serves as input into a LLM.
-(Zhao et al., 2024; Wan et al., 2024; Tang et al., 2024b; Choi et al., 2023).
-In order to get good latent representations of the ECGs, an encoder is first pretrained on a self-supervised learning (SSL) objective (e.g., contrastive learning, masked language/image modeling).
+Previous works have largely focused on representing ECG data by feeding the raw signal into a neural network encoder to output a latent representation, which is then used as input for an LLM (Zhao et al., 2024; Wan et al., 2024; Tang et al., 2024b).
+Recent studies have explored using printed ECGs as images for natural language generation (NLG) (Liu et al., 2024c; Khunte et al., 2024).
+However, our study focuses solely on methods that utilize the direct ECG signal.
+In order to obtain robust latent representations of the ECG signal, an encoder is first trained on a self-supervised learning (SSL) objective (e.g., contrastive learning, masked language/image modeling).
 Although model performance in terms of label classification has been excellent when using these approaches, we want to be able to generate soft labels akin to clinical notation since ECGs often have overlapping and non-mutually exclusive descriptors.
-There have been some efforts in this direction (Qiu et al., 2023a; Tang et al., 2024b; Wan et al., 2024; Fu et al., 2024; Zhao et al., 2024), however, they suffer in efficiency (i.e., requires 2 stages of training).
-In our work, we challenge this 2-stage pretraining approach by transforming the ECG into tokens using ECG-Byte and directly training a LLM for NLG.
+While some efforts have been made in this direction (Tang et al., 2024b; Zhao et al., 2024), they face challenges in efficiency (requiring two stages of training) and interpretability (as the latent feature vectors from the ECG encoder are not interpretable).
+In our work, we challenge these two-stage pretraining approaches by transforming the ECG into tokens using ECG-Byte and directly training an LLM for NLG.
 2.3 Byte Pair Encoding for Domains Outside of Language
-The Byte Pair Encoding (BPE) algorithm was first introduced by Gage (1994) for data compression.
-It was later adapted to the natural language processing (NLP) domain (Sennrich et al., 2016) and has been the favored approach to tokenization for the most popular language models (Grattafiori et al., 2024; Brown et al., 2020) due to its efficiency and robustness to rare words.
-Byte pair encoding has seen success in representing modalities outside of language, including molecular graphs (Shen and Póczos, 2024), electroencephalogram (EEG) (Klymenko et al., 2023), and more generally, physiological signals.(Tavabi and Lerman, 2021).
-However, in theses cases the byte pair encoded representations are simply used for classification.
+The Byte Pair Encoding (BPE) algorithm was introduced by Gage (1994) for data compression and later adapted for natural language processing (NLP) (Sennrich et al., 2016).
+It is favored for tokenization in popular language models (Grattafiori et al., 2024; Brown et al., 2020) due to its efficiency and robustness to rare words.
+Beyond language, BPE has been applied to modalities such as molecular graphs (Shen and Póczos, 2024), electroencephalograms (EEG) (Klymenko et al., 2023), and other physiological signals (Tavabi and Lerman, 2021) for classification.
 Most recently, Tahery et al.
-(2024) leveraged quantization and BPE to compress ECG signals and pass as inputs to a BERT (Devlin et al., 2019) model for SSL.
-However, in their work, they only use this representation for classification.
-As previously mentioned, we believe classification alone may limit aspects of ECG interpretation, thus we utilize these representations for generative diagnosis.
-Additionally, previous works utilized a pre-existing BPE tokenizer based on SentencePiece (Kudo and Richardson, 2018), and do not conduct further analysis of how the BPE algorithm is merging the ECGs.
+(2024) used quantization and BPE to compress ECG signals as inputs to a BERT (Devlin et al., 2019) model for self-supervised learning, though only for classification.
+We argue that classification alone may limit ECG interpretation and therefore leverage these representations for generative diagnosis.
+Additionally, previous works employ a pre-existing BPE tokenizer (Tahery et al., 2024) based on SentencePiece (Kudo and Richardson, 2018) without further analyzing how the BPE algorithm merges ECGs.
+Inspired by HuggingFace (Wolf et al., 2020), we develop ECG-Byte, a custom Rust-based BPE tokenizer for ECGs.
+We conduct an extensive analysis of merged token usage, the distribution of tokenized ECG lengths, and provide visualizations mapping compressed tokens to the original ECG signal.
 3 Methods
@@
 The baselines we compare our results with all utilize the single-verify, single-choose, and single-query categorized questions from the ECG-QA dataset.
-single-verify corresponds to yes or no questions, single-choose corresponds to where a selected answer is made from two given options, and single-query consists of open-ended questions.
-The ECG signals collected from both both datasets (i.e., MIMIC-IV ECG and PTB-XL) are sampled at 500 Hz for 10 seconds, resulting in a 5000 length, 12 lead ECG.
+The ECG signals collected from both datasets (i.e., MIMIC-IV ECG and PTB-XL) are sampled at 500 Hz for 10 seconds, resulting in a 5000 length, 12 lead ECG.
 Preprocessing
-We preprocess all datasets used in the study in the same manner to maintain consistency.
-We first convert the ordering of the leads for the MIMIC-IV ECG dataset (i.e., [‘I’, ‘II’, ‘III’, ‘aVR’, ‘aVF’, ‘aVL’, ‘V1’, ‘V2’, ‘V3’, ‘V4’, ‘V5’, ‘V6’])to the PTB-XL dataset format (i.e., [‘I’, ‘II’, ‘III’, ‘aVL’, ‘aVR’, ‘aVF’, ‘V1’, ‘V2’, ‘V3’, ‘V4’, ‘V5’, ‘V6’]).
-We then use a notch filter at 50 Hz and 60 Hz to remove powerline interference.
-Each frequency is targeted with a quality factor of 30 to minimize distortion, and filtering is applied bidirectionally to prevent phase shifts.
-Next, a fourth-order Butterworth bandpass filter with a range of 0.5–100 Hz isolates relevant ECG components while attenuating high-frequency noise and low-frequency drift.
-To address baseline wander caused by respiratory or movement artifacts, we bidirectionally apply a fourth-order Butterworth highpass filter with a cutoff of 0.05 Hz.
-After filtering, we apply wavelet denoising to further reduce noise.
-Using the Daubechies-6 (db6) wavelet at level 4, we decompose each ECG signal into wavelet coefficients.
-A soft threshold, based on the median absolute deviation of the detail coefficients, is applied to each coefficient level to suppress noise, ensuring values near zero are excluded from reconstruction.
-Since 250 Hz is a generally accepted sampling frequency adequate for heartbeat analysis (Kwon et al., 2018), we downsample the 500 Hz sampling frequency to 250.
-We then segment the 10 second signal to non-overlapping windows of 2 seconds, and use each 12 lead 2 second segment of the ECG signal as input to the model.
-However, for training the tokenizer, we did not want to introduce this discontinuity across the full 10 seconds.
-Thus, we utilize the unsegmented, 10 second ECG signal for training ECG-Byte.
-Lastly, during the unsegmented preprocessing pipeline, we record the global 1st and 99th percentiles out of 300,000 samples to utilize in our later steps of training ECG-Byte for normalization.
+All datasets are preprocessed uniformly.
+For the MIMIC-IV ECG, we first reorder leads to match the PTB-XL format (from [I, II, III, aVR, aVF, aVL, V1–V6] to [I, II, III, aVL, aVR, aVF, V1–V6]).
+Powerline interference is removed using bidirectional notch filters at 50 Hz and 60 Hz (Q=30).
+A fourth-order Butterworth bandpass filter (0.5–100 Hz) isolates ECG components, while a bidirectional fourth-order highpass filter (cutoff 0.05 Hz) mitigates baseline wander.
+We then apply Daubechies-6 (db6) wavelet denoising at level 4, using a soft threshold based on the median absolute deviation of the detail coefficients.
+The signal is downsampled from 500 Hz to 250 Hz and segmented into non-overlapping 2-second windows for model input—except during tokenizer training, where the full 10-second signal is retained to avoid discontinuity.
+Finally, global 1st and 99th percentiles (from 300,000 samples) are recorded for later normalization in training ECG-Byte.
 3.2 ECG as Bytes
@@
 If K-means fails to yield distinct clusters, DBSCAN (Ester et al., 1996) is used as a fallback.
-Stratified sampling is performed by randomly selecting ECGs from each cluster in proportion to its size, resulting in a total sample of 200,000 ECGs.
+Stratified sampling is performed by randomly selecting ECGs from each cluster in proportion to its size, resulting in a total sample of 200,000 ECGs for training ECG-Byte.
 Quantization
@@
 This transformation shifts and scales $X$ so that the normalized values fall within the range [0, 1].
-We then apply clipping to ensure that values remain strictly within this range:
-$X_{\text{clipped}}=\text{clip}(X_{\text{norm}},0,1)$ (2)
-Inspired by previous works (Klymenko et al., 2023; Tavabi and Lerman, 2021; Chen et al., 2022; Han et al., 2024), we quantize $X_{\text{clipped}}$ into discrete levels for symbolic representation.
+We then apply clipping to ensure that values remain strictly within this range.
+Inspired by previous works (Klymenko et al., 2023; Tavabi and Lerman, 2021; Chen et al., 2022; Han et al., 2024), we quantize $X_{\text{norm}}$ into discrete levels for a symbolic representation.
 Let $\mathcal{A}$ be the set of 26 symbols, corresponding to the lowercased letters in the English alphabet, $\mathcal{A}=\{\text{a},\text{b},\ldots,\text{z}\}$.
 The alphabet size $|\mathcal{A}|=26$ defines the number of discrete levels.
-We scale and floor $X_{\text{clipped}}$ to integer values, then take the minimum between the floored value and the maximum number of bins as the following:
-$X_{\text{quant}}=\text{min}(\lfloor X_{\text{clipped}}\times|\mathcal{A}|\rfloor,(|\mathcal{A}|-1))$ (3)
+We scale and floor $X_{\text{norm}}$ to integer values, then take the minimum between the floored value and the maximum number of bins as the following:
+$X_{\text{quant}}=\text{min}(\lfloor X_{\text{norm}}\times|\mathcal{A}|\rfloor,(|\mathcal{A}|-1))$ (2)
 Finally, each integer value in $X_{\text{quant}}$ is mapped to a corresponding symbol in $\mathcal{A}$ to yield the symbolic signal, which serves as a discrete representation of the ECG.
-After transforming each ECG signal instance into its symbolic form, we first flatten each symbolic ECG instance $X_{\text{quant}}^{(i)}$ into a 1-dimensional sequence of symbols:
-$X_{\text{symb}}^{(i)}=\text{flatten}(X_{\text{quant}}^{(i)}),\quad X_{\text{symb}}^{(i)}\in\mathcal{A}^{CT}$ (4)
-where $i$ indexes over all instances in the dataset, and $X_{\text{symb}}^{(i)}$ is the flattened sequence of symbols of length $C\cdot T$.
-Next, we concatenate all flattened instances $X_{\text{symb}}^{(1)},X_{\text{symb}}^{(2)},\ldots,X_{\text{symb}}^{(N)}$ across the entire dataset to form a single, long symbolic sequence:
-$\mathbf{X}_{\text{concat}}=X_{\text{symb}}^{(1)}\|X_{\text{symb}}^{(2)}\|\cdots\|X_{\text{symb}}^{(N)},\quad\mathbf{X}_{\text{concat}}\in\mathcal{A}^{NCT}$ (5)
-[…]
+After transforming each ECG signal instance into its symbolic form, we first flatten each symbolic ECG instance $X_{\text{quant}}^{(i)}$ into a 1-dimensional sequence of symbols $X_{\text{symb}}^{(i)}$, where $X_{\text{symb}}^{(i)}\in\mathcal{A}^{CT}$, $i$ indexes over all instances in the dataset, and $X_{\text{symb}}^{(i)}$ is the flattened sequence of symbols of length $C\cdot T$.
+Next, we concatenate all flattened instances $X_{\text{symb}}^{(1)},X_{\text{symb}}^{(2)},\ldots,X_{\text{symb}}^{(N)}$ across the entire dataset to form a single, long symbolic sequence $X_{\text{concat}}$, where $X_{\text{concat}}\in\mathcal{A}^{NCT}$, and $N$ is the total number of instances in the dataset.
+The concatenated symbolic sequence $X_{\text{concat}}$ of length $N\cdot C\cdot T$ is then used to train ECG-Byte.
 ECG-Byte Training Process
-After obtaining the string representation $\mathbf{X}_{\text{concat}}$ of the ECG dataset, we train ECG-Byte to compress the discretized ECG signals by iteratively merging the most frequent byte pairs into single tokens, following the BPE algorithm.
-The process starts by converting $\mathbf{X}_{\text{concat}}$ into an ID vector of 8-bit unsigned integers and initializing a vocabulary map (vocab) for string representations of bytes and a vocab_tokens map to encode bytes as singleton lists.
+After obtaining the string representation $\mathbf{X}_{\text{concat}}$ of the ECG dataset, we train ECG-Byte to compress the discretized ECG signals by iteratively merging the most frequent byte pairs into single tokens, following the BPE algorithm (Gage, 1994).
+The process starts by converting $\mathbf{X}_{\text{concat}}$ into a vector of token IDs derived from 8-bit byte values (stored as 32-bit unsigned integers) and initializing a vocabulary map (vocab) for string representations of bytes and a vocab_tokens map to encode bytes as singleton lists.
 IDs and vocab are initialized to cover the full byte range (0–255), mapping symbols in $\mathcal{A}$ to ASCII values (97–122), while reserving other byte values for unknown bytes.
@@
 For each merge iteration, ECG-Byte calculates adjacent byte pair frequencies using a parallelized get_stats function, efficiently aggregating counts via a fold-and-reduce strategy.
-The most frequent pair is identified as the ”best pair” to merge, and the merge function replaces occurrences of this pair in the ID vector with a new token ID, extending the vocabulary and updating vocab_tokens accordingly.
+The most frequent pair is identified as the “best pair” to merge, and the merge function replaces occurrences of this pair in the ID vector with a new token ID, extending the vocabulary and updating vocab_tokens accordingly.
 This process repeats until the specified number of merges is reached or no pairs remain.
@@
 ECG-Byte, implemented in Rust for speed, provides a lightweight, flexible framework for representing ECG signals as discrete tokens while drawing inspiration from HuggingFace’s tokenizer (Wolf et al., 2020).
-Detailed pseudocode for the main training pipeline is provided in Algorithm 1, with merge and get_stats functions detailed in Appendix A.1.
+We provide the detailed pseudocode of the training process in Algorithm 1 and pseudocode of the merge and get_stats functions in Appendix B.1.
+Algorithm 1 Training Process for ECG-Byte
+Input: Input $X_{\text{concat}}$, number of merges num_merges.
+Output: Tuple containing final encoded IDs, vocabulary map, and merge history.
+1:  ids $\leftarrow$ Convert $X_{\text{concat}}$ to a vector of $u32$
+2:  vocab $\leftarrow$ Mappings from IDs 00 to $255$ as string
+3:  vocab_tokens $\leftarrow$ Mappings from IDs 00 to $255$ to singleton lists
+4:  merges $\leftarrow$ Empty list
+5:  for $i\leftarrow 0$ to num_merges - 1 do
+6:     pairs $\leftarrow$ get_stats(ids)
+7:     if pairs is empty then
+8:        break
+9:     end if
+10:     best_pair $\leftarrow$ pair in pairs with highest frequency
+11:     if best_pair is not found then
+12:        break
+13:     end if
+14:     new_id $\leftarrow 256+i$
+15:     ids $\leftarrow$ merge(ids, best_pair, new_id)
+16:     vocab[new_id] $\leftarrow$ concat(vocab[best_pair.0], vocab[best_pair.1])
+17:     new_token $\leftarrow$ concat(vocab_tokens[best_pair.0], vocab_tokens[best_pair.1])
+18:     vocab_tokens[new_id] $\leftarrow$ new_token
+19:     Append (new_token, new_id) to merges
+20:  end for
+21:  return (ids, vocab, merges)
 ECG-Byte Encoding Process
-After training ECG-Byte, we encode any quantized ECG signal $X_{\text{symb}}$ by first converting each byte in the ECG signal to a 32-bit unsigned integer and building a trie structure, where each node represents a byte or a merged token sequence from prior encoding steps.
+After training ECG-Byte, we encode any quantized ECG signal $X_{\text{symb}}$ by first converting each byte to a 32-bit unsigned integer and building a trie structure, where each node represents a byte or a merged token sequence from prior encoding steps.
 The trie is initialized with single-byte tokens (0-255) and is extended with custom token sequences from the learned merge history.
@@
 The final encoded sequence is returned as output_ids, where we will denote as $X_{\text{ID}}$.
-We provide the detailed pseudocode of the encoding process in Algorithm 2.
-Input: Input $X_{concat}$, Number of merges num_merges
-Output: Tuple containing final encoded IDs, vocabulary map, and merge history
-Function byte_pair_encoding($X_{concat}$, num_merges):
-       Convert $X_{concat}$ to ID vector ids by casting each byte to $u32$;
-       Initialize vocab with mappings from IDs 00 to $255$ to their string representations;
-       Initialize vocab_tokens with mappings from IDs 00 to $255$ to singleton lists;
-       Initialize empty list merges;
-      for $i\leftarrow 0$ to num_merges exclusive do
-             pairs $\leftarrow$ get_stats(ids) using parallel processing;
-            if pairs is empty then
-                   break
-             end if
-            best_pair $\leftarrow$ Pair in pairs with highest frequency;
-             if best_pair is not found then
-                   break
-             end if
-            new_id $\leftarrow 256+i$;
-             ids $\leftarrow$ merge(ids, best_pair, new_id);
-            vocab[new_id] $\leftarrow$ concat(vocab[best_pair.0], vocab[best_pair.1]);
-            new_token $\leftarrow$ concat(vocab_tokens[best_pair.0], vocab_tokens[best_pair.1]);
-             vocab_tokens[new_id] $\leftarrow$ new_token;
-            merges.append(new_token, new_id);
-       end for
-      return (ids, vocab, merges);
-Algorithm 1 Training Process for ECG-Byte
-Input: Input $X_{\text{symb}}$, Merge history merges containing pairs of token sequences and their token IDs
-Output: Vector of encoded IDs
-Function encode($X_{\text{symb}}$, merges):
-       ids $\leftarrow$ Convert $X_{\text{symb}}$ to vector of u32 by casting each byte;
-       Initialize root TrieNode trie_root using TrieNode::new();
-      for b $\leftarrow 0$ to $255$ do
-             insert(trie_root, [b], b);
-       end for
-      foreach (token_sequence, token_id) in merges do
-             insert(trie_root, token_sequence, token_id);
-       end foreach
-      Initialize empty list output_ids;
-       i $\leftarrow 0$;
-      while i $<$ length of ids do
-             node $\leftarrow$ trie_root;
-             match_len $\leftarrow 0$;
-             match_id $\leftarrow$ None;
-            for j $\leftarrow$ i to length of ids do
-                   id $\leftarrow$ ids[j];
-                  if id exists in node.children then
-                         node $\leftarrow$ node.children[id];
-                        if node.token_id is not None then
-                               match_len $\leftarrow$ j - i + 1;
-                               match_id $\leftarrow$ node.token_id;
-                         end if
-                   end if
-                  else
-                         break
-                   end if
-             end for
-            if match_id is not None then
-                   output_ids.append(match_id);
-                   i $\leftarrow$ i + match_len;
-             end if
-            else
-                   output_ids.append(ids[i]);
-                   i $\leftarrow$ i + 1;
-             end if
-       end while
-      return output_ids;
-Algorithm 2 Encoding Process for ECG-Byte
+We provide the detailed pseudocode of the encoding process in Algorithm 4 in the Appendix due to page limitations.
 3.3 Large Language Model
 In this study, we utilize the Llama-3.2-1B (Grattafiori et al., 2024) checkpoint through the HuggingFace API (Wolf et al., 2020) unless specified otherwise.
-The Llama 3.2 series model is a variant of the Llama 3 models (Grattafiori et al., 2024) and support context lengths of up to 128k tokens.
-They are notable for their superior performance despite having 1 billion parameters, making them highly efficient and capable models to test our methodology upon.
 We also provide an ablation study in subsection 5.4 where we utilize other popular LLMs, such as GPT2 XL 1.5B (Radford et al., 2019), Gemma 2B (Team et al., 2024), and OPT 1.3B (Zhang et al., 2022).
+With the exception of modifying the token embedding size (since we add new ECG tokens), we utilize all default hyperparameters out-of-the-box unless specified otherwise.
 3.4 Learning Objective
-The learning objective for training the LLM considers a sequence composed of three parts, $\{X_{\text{ID}},Q,\mathcal{S}\}$, where $X_{\text{ID}}\in\mathcal{V}^{M}$ represents the encoded ECG sequence of length $l_{X_{\text{ID}}}=|X_{\text{ID}}|$, with each token drawn from the extended vocabulary $\mathcal{V}$ of size $M$, $Q$ represents the tokenized question, and $\mathcal{S}$ denotes the tokenized answer sequence.
-The input sequence includes special tokens: - [BOS] as the beginning-of-sequence token, - [SIG_START] and [SIG_END] to indicate the start and end of the encoded ECG sequence, and - [EOS] as the end-of-sequence token for the generated answer.
-The motivation for adding [SIG_START] and [SIG_END] special tokens is inspired by Liu et al.
-(2023), where they utilize special tokens indicating the start and end of the image.
-Thus, the full input sequence is structured as:
-[…]
-[…]
-$L=1+1+l_{X_{\text{ID}}}+1+l_{Q}+l_{\mathcal{S}}+1,$
 […]
@@
 4.1 Experimental Settings
-We fine-tuned the LLM using the AdamW optimizer (Kingma and Ba, 2017) with a learning rate of $41e-$ 4, weight decay of $21e-$ 2, and a custom learning rate scheduler.
-This scheduler applies an initial learning rate $\mathrm{init\_lr}$ scaled by the model’s hidden dimension ($\mathrm{d_{model}}^{-0.5}$) and dynamically adjusts it based on training steps, with a warm-up phase of 500 steps.
-The learning rate at step $\mathrm{n_{steps}}$ is updated as $\text{lr}=\mathrm{init\_lr}\times\min\left(\mathrm{n_{steps}}^{-0.5},\ \mathrm{n_{warmup}}^{-1.5}\times\mathrm{n_{steps}}\right)$.
-We set $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1e-8$, batch size 2, and trained for 1 epoch.
-Additionally, we only train on a randomly sampled subset of 400,000 ECG instances for each respective dataset due to computational resources, unless specified otherwise.
+We finetuned the LLM using the AdamW optimizer (Kingma and Ba, 2017) with a learning rate of $41e-$ 4, weight decay of $21e-$ 2 (Krogh and Hertz, 1991), and a custom learning rate scheduler.
+We provide details about the learning rate scheduler in Appendix A.1 due to page limitations.
+We train the LLMs for 1 epoch and with a batch size of 2.
+We set the exponential decay rates to be $\beta_{1}=0.9$ and $\beta_{2}=0.99$.
+We set $\epsilon=1e-8$ as a constant for numerical stability.
+Due to computational constraints, we limit each dataset to a randomly sampled training subset of 400,000 ECG instances and an independent inference subset of 25,000 instances, unless specified otherwise.
 We also utilize LoRA (Hu et al., 2021) to finetune the LLM with rank = 16, $\alpha_{LoRA}=32$, and dropout = 0.05.
@@
 We use popular metrics for NLG namely the BLEU-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Meteor (Banerjee and Lavie, 2005), and BertScore F1 (Zhang et al., 2020) metrics.
-[…]
-tab:main_results Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1 ECG-Chat (Zhao et al., 2024) MIMIC-IV ECG Pretrain PTB-XL 11.19 29.93 35.10 - $L_{CL}$ 8.10 ± 0.25 31.36 ± 0.31 27.55 ± 0.36 89.35 ± 0.04 $L_{MIM}$ 6.21 ± 0.22 30.63 ± 0.13 24.91 ± 0.14 90.44 ± 0.04 $L_{MERL}$ (Liu et al., 2024) 10.22 ± 0.25 32.95 ± 0.12 25.60 ± 0.17 89.94 ± 0.01 $L_{CL}$ + $L_{MIM}$ 9.33 ± 0.22 30.45 ± 0.21 24.37 ± 0.36 90.29 ± 0.02 ECG-Byte 11.00 ± 0.19 33.41 ± 0.05 24.95 ± 0.09 90.02 ± 0.01 $L_{CL}$ ECG-QA MIMIC-IV ECG-QA MIMIC-IV 10.22 ± 0.06 38.41 ± 0.48 24.66 ± 0.23 90.42 ± 0.09 $L_{MIM}$ 7.90 ± 0.23 29.28 ± 0.38 19.03 ± 0.11 67.91 ± 0.17 $L_{MERL}$ (Liu et al., 2024) 10.95 ± 0.24 38.18 ± 0.58 26.24 ± 0.36 90.80 ± 0.06 $L_{CL}+L_{MIM}$ 8.57 ± 0.14 34.00 ± 0.25 25.22 ± 0.30 87.72 ± 0.04 ECG-Byte 11.23 ± 0.12 42.49 ± 0.53 27.08 ± 0.15 91.30 ± 0.04 $L_{CL}$ ECG-QA PTB-XL ECG-QA PTB-XL 8.89 ± 0.25 28.63 ± 0.47 18.45 ± 0.31 72.63 ± 0.40 $L_{MIM}$ 15.14 ± 0.28 46.71 ± 0.41 29.64 ± 0.30 92.12 ± 0.10 $L_{MERL}$ (Liu et al., 2024) 13.84 ± 0.19 40.14 ± 0.39 26.24 ± 0.35 91.88 ± 0.09 $L_{CL}+L_{MIM}$ 14.72 ± 0.27 42.88 ± 0.13 28.25 ± 0.27 89.40 ± 0.01 ECG-Byte 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
+All reported NLG results are presented as means with standard deviations computed over 5 random seeds.
 Table 1: NLG mean results with standard deviations over 5 random seeds comparing against different baselines.
+Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1
+ECG-Chat (Zhao et al., 2024) MIMIC-IV ECG Pretrain PTB-XL 11.19 29.93 35.10 -
+$L_{CL}$ 8.10 ± 0.25 31.36 ± 0.31 27.55 ± 0.36 89.35 ± 0.04
+$L_{MIM}$ 6.21 ± 0.22 30.63 ± 0.13 24.91 ± 0.14 90.44 ± 0.04
+$L_{MERL}$ (Liu et al., 2024a) 10.22 ± 0.25 32.95 ± 0.12 25.60 ± 0.17 89.94 ± 0.01
+$L_{Dual}$ 9.33 ± 0.22 30.45 ± 0.21 24.37 ± 0.36 90.29 ± 0.02
+ECG-Byte 11.00 ± 0.19 33.41 ± 0.05 24.95 ± 0.09 90.02 ± 0.01
+$L_{CL}$ ECG-QA MIMIC-IV ECG-QA MIMIC-IV 10.22 ± 0.06 38.41 ± 0.48 24.66 ± 0.23 90.42 ± 0.09
+$L_{MIM}$ 7.90 ± 0.23 29.28 ± 0.38 19.03 ± 0.11 67.91 ± 0.17
+$L_{MERL}$ (Liu et al., 2024a) 10.95 ± 0.24 38.18 ± 0.58 26.24 ± 0.36 90.80 ± 0.06
+$L_{Dual}$ 8.57 ± 0.14 34.00 ± 0.25 25.22 ± 0.30 87.72 ± 0.04
+ECG-Byte 11.23 ± 0.12 42.49 ± 0.53 27.08 ± 0.15 91.30 ± 0.04
+$L_{CL}$ ECG-QA PTB-XL ECG-QA PTB-XL 8.89 ± 0.25 28.63 ± 0.47 18.45 ± 0.31 72.63 ± 0.40
+$L_{MIM}$ 15.14 ± 0.28 46.71 ± 0.41 29.64 ± 0.30 92.12 ± 0.10
+$L_{MERL}$ (Liu et al., 2024a) 13.84 ± 0.19 40.14 ± 0.39 26.24 ± 0.35 91.88 ± 0.09
+$L_{Dual}$ 14.72 ± 0.27 42.88 ± 0.13 28.25 ± 0.27 89.40 ± 0.01
+ECG-Byte 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
 5 Results
 5.1 Natural Language Generation
-We present our main results in Table LABEL:tab:main_results, comparing ECG-Byte with prior works and self-implemented two-stage pretraining methods.
+We present our main results in Table 1, comparing ECG-Byte with prior works and self-implemented two-stage pretraining methods.
 Notably, Zhao et al.
-(2024) is not directly comparable due to differing data splits and pretraining datasets, though reported metrics use the same datasets.
+(2024) is not directly comparable due to differing data splits and pretraining datasets.
 Zhao et al.
-(2024) train on the full MIMIC-IV ECG Pretrain dataset, finetune on an instruction-tuning dataset for ECG-related conversations, and evaluate on PTB-XL (Wagner et al., 2020) using a unified question: “Could you please help me explain my ECG?” To establish baselines, we implement generic two-stage pretraining methods: $L_{CL}$, $L_{MIM}$, and $L_{CL}+L_{MIM}$.
-Here, $L_{CL}$ employs contrastive learning (Liu et al., 2024; Gopal et al., 2021; Pham et al., 2024; Kiyasseh et al., 2021), $L_{MIM}$ uses Masked Image Modeling (MIM) (Choi et al., 2023; Na et al., 2024; Yang et al., 2022), and $L_{CL}+L_{MIM}$ combines both (Oh et al., 2022; McKeen et al., 2024).
-These approaches utilize pretrained CLIP (Radford et al., 2021) and ViT (Dosovitskiy et al., 2021), where ECG signals are transformed into three-channel images for finetuning.
-We adapt Liu et al.
-(2024)’s state-of-the-art contrastive method ($L_{MERL}$) for fair comparison.
-Their most effective model uses a 1D ResNet backbone (He et al., 2015); hence, we employ ResNet101 for direct ECG signal processing.
-For $L_{CL}$, $L_{MIM}$, $L_{CL}+L_{MIM}$, and $L_{MERL}$, training is conducted on the full, preprocessed MIMIC-IV ECG dataset with a batch size of 64 during the first stage.
-Implementation details for both training stages are in Appendix B.
-Table LABEL:tab:main_results demonstrates ECG-Byte’s effectiveness, showing competitive or superior performance across all metrics and datasets compared to other methods.
-Qualitative examples are provided in Appendix C.2.
-5.2 Cross Dataset Transferability
-We present the results of cross-dataset transferability in Table LABEL:tab:cross, comparing our approach, ECG-Byte, with two-stage pretraining methods.
-ECG-Byte achieves the best zero-shot transfer performance from the ECG-QA PTB-XL dataset to the ECG-QA MIMIC-IV dataset.
-When transferring from the ECG-QA MIMIC-IV dataset to the ECG-QA PTB-XL dataset, although other 2-stage pretraining methods demonstrate higher performance, ECG-Byte maintains competitive results across all metrics.
+(2024) train on the full MIMIC-IV ECG Pretrain dataset, finetune on an instruction-tuning dataset for ECG-related conversations, and evaluate on PTB-XL (Wagner et al., 2020) using a unified question: “Could you please help me explain my ECG?” We also note that ECG-Chat (Zhao et al., 2024) employs numerous generation enhancements such as Retrieval Augmented Generation (RAG), DSPy for automatic prompt tuning, and an adapted Diagnosis-Driven Prompt (DDP) inspired by Jin et al.
+(2024).
+In our work, we only observe performance of NLG given a text prompt and ECG signal.
+Therefore, to establish comparable baselines, we implement generic two-stage pretraining methods utilizing strong encoders: $L_{CL}$, $L_{MIM}$, and $L_{Dual}=L_{CL}+L_{MIM}$.
+Here, $L_{CL}$ employs contrastive learning (Liu et al., 2024a; Gopal et al., 2021; Pham et al., 2024; Kiyasseh et al., 2021), $L_{MIM}$ uses Masked Image Modeling (MIM) (Choi et al., 2023; Na et al., 2024; Yang et al., 2022), and $L_{Dual}$ combines both (Oh et al., 2022; McKeen et al., 2024).
+All three implementations utilize pretrained CLIP (Radford et al., 2021) and ViT (Dosovitskiy et al., 2021) models, where ECG signals are transformed into three-channel images for finetuning.
+Additionally, we adapt Liu et al.
+(2024a)’s state-of-the-art contrastive method ($L_{MERL}$) for fair comparison.
+Their most effective model uses a 1D ResNet backbone (He et al., 2015), therefore we utilize the ResNet101 variant.
+For $L_{CL}$, $L_{MIM}$, $L_{Dual}$, and $L_{MERL}$, training is conducted on the full, preprocessed MIMIC-IV ECG dataset with a batch size of 64 during the first stage to simulate prior work settings.
+Implementation details for both training stages are in Appendix C.
+Table 1 demonstrates ECG-Byte’s effectiveness, showing competitive or superior performances across all metrics and datasets compared to other methods.
+We also highlight that ECG-Byte was initially trained on a separate, unsegmented subset of 200,000 ECGs from the MIMIC-IV ECG dataset.
+With this fact, ECG-Byte displays strong performances when training and inferencing the LLM on ECGs of different lengths (i.e., segmented to 2 seconds) and from another dataset (i.e., PTB-XL).
+Qualitative examples of generated answers are provided in Appendix D.4.
 5.2 Efficiency of ECG-Byte
-We compare the efficiency of our end-to-end approach using ECG-Byte with 2-stage pretraining methods in Table LABEL:tab:efficiency.
-First, we examine the amount of data required for each method.
-As previously noted, the first stage of the two-stage pretraining methods is trained on the full MIMIC-IV ECG dataset (Johnson et al., 2023) using segmented ECGs, which are also used as input during the second stage.
-While ECG-Byte is trained on unsegmented ECGs, we convert the number of unsegmented ECGs to an equivalent count of segmented ECGs.
-Additionally, the reduced data requirement for ECG-Byte is due to our sampling approach, where only a subset of ECGs is used to train the tokenizer.
-Under these settings, our proposed method achieves competitive results using approximately $\sim$48of the data required for 2-stage pretraining methods.
-In terms of training time, our approach requires less than half the time needed for two-stage pretraining.
-The training time for the two-stage methods is averaged across our self-implemented approaches ($L_{CL}$, $L_{MIM}$, and $L_{CL}+L_{MIM}$) and the $L_{MERL}$ method proposed by Liu et al.
-(2024).
-[…]
-tab:cross Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1 $L_{CL}$ ECG-QA MIMIC-IV ECG-QA PTB-XL 11.64 ± 0.45 41.48 ± 0.11 25.74 ± 0.13 91.24 ± 0.05 $L_{MIM}$ 11.70 ± 0.29 42.22 ± 0.28 26.41 ± 0.10 91.51 ± 0.03 $L_{MERL}$ (Liu et al., 2024) 11.53 ± 0.19 39.23 ± 0.40 25.58 ± 0.28 91.59 ± 0.03 $L_{CL}$ + $L_{MIM}$ 9.71 ± 0.10 35.10 ± 0.28 24.91 ± 0.19 87.88 ± 0.08 ECG-Byte 8.70 ± 0.04 40.39 ± 0.40 23.29 ± 0.18 91.51 ± 0.03 $L_{CL}$ ECG-QA PTB-XL ECG-QA MIMIC-IV 5.10 ± 0.04 22.77 ± 0.28 14.63 ± 0.32 77.89 ± 0.13 $L_{MIM}$ 7.68 ± 0.46 35.77 ± 0.13 22.32 ± 0.33 90.28 ± 0.07 $L_{MERL}$ (Liu et al., 2024) 7.39 ± 0.15 28.33 ± 0.58 18.59 ± 0.35 89.30 ± 0.05 $L_{CL}$ + $L_{MIM}$ 7.49 ± 0.21 30.53 ± 0.59 20.25 ± 0.27 86.53 ± 0.11 ECG-Byte 7.86 ± 0.13 35.01 ± 0.41 21.49 ± 0.24 90.29 ± 0.07
-Table 2: Mean results with standard deviations over 5 random seeds on zero shot cross-dataset transferability.
-[…]
-tab:efficiency 1st Stage 2nd Stage ECG-Byte end-to-end # of Data 2,513,435 400,000 1,000,000 400,000 Total # of Data 2,913,435 1,400,000 Time (minutes) $\sim$1258.50 $\sim$469.25 $\sim$385.12 $\sim$420.32 Total Time (minutes) $\sim$1727.75 $\sim$805.44
-Table 3: Efficiency of our method compared against 2-stage pretraining methods.
+We compare the efficiency of our end-to-end (end-to-end A/B) approach (ECG-Byte A/B) with two-stage pretraining (1st Stage / 2nd Stage) in Table 2.
+The top and middle sections report data and time requirements for both settings used in our study.
+The two-stage methods first train on the full MIMIC-IV ECG dataset (Johnson et al., 2023) using segmented ECGs, which are then used as inputs in the second stage.
+Although ECG-Byte is trained on unsegmented ECGs, we convert these counts to segmented-equivalents and reduce data requirements by sampling only a subset for tokenizer training.
+Training time for two-stage methods is averaged over our self-implemented approaches ($L_{CL}$, $L_{MIM}$, $L_{Dual}$) and the $L_{MERL}$ method (Liu et al., 2024a).
+Under these settings, our method achieves competitive results using approximately 48% of the data and is about 3 times faster in total training time.
+Even when training ECG-Byte on 500,000 unsegmented ECGs (equivalent to 2,500,000 segmented ECGs) as seen in the bottom section (ECG-Byte B), our method remains 2.38 times faster than two-stage pretraining, highlighting its efficiency.
+Table 2: Efficiency of our method compared against two-stage pretraining methods.
+Method # of Data Total # of Data Time (min.) Total Time (min.)
+1st Stage 2,513,435 2,913,435 $\sim$1258.50 $\sim$1727.75
+2nd Stage 400,000 $\sim$469.25
+ECG-Byte A 1,000,000 1,400,000 $\sim$152.47 $\sim$572.79
+end-to-end A 400,000 $\sim$420.32
+ECG-Byte B 2,500,000 2,900,000 $\sim$304.27 $\sim$724.59
+end-to-end B 400,000 $\sim$420.32
+Table 3: Mean results with standard deviations over 5 random seeds on training two-stage pretraining methods in an end-to-end fashion.
+Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1
+$L_{MIM}$ ECG-QA MIMIC-IV ECG-QA MIMIC-IV 7.90 ± 0.23 29.28 ± 0.38 19.03 ± 0.11 67.91 ± 0.17
+End-to-End $L_{MIM}$ 7.71 ± 0.36 28.19 ± 0.13 18.88 ± 0.54 67.82 ± 0.04
+$L_{MERL}$ (Liu et al., 2024a) 10.95 ± 0.24 38.18 ± 0.58 26.24 ± 0.36 90.80 ± 0.06
+End-to-End $L_{MERL}$ (Liu et al., 2024a) 8.26 ± 0.17 34.88 ± 0.26 22.14 ± 0.33 86.64 ± 0.03
+ECG-Byte 11.23 ± 0.12 42.49 ± 0.53 27.08 ± 0.15 91.30 ± 0.04
+$L_{MIM}$ ECG-QA PTB-XL ECG-QA PTB-XL 15.14 ± 0.28 46.71 ± 0.41 29.64 ± 0.30 92.12 ± 0.10
+End-to-End $L_{MIM}$ 12.04 ± 0.17 40.44 ± 0.25 24.84 ± 0.43 87.71 ± 0.11
+$L_{MERL}$ (Liu et al., 2024a) 13.84 ± 0.19 40.14 ± 0.39 26.24 ± 0.35 91.88 ± 0.09
+End-to-End $L_{MERL}$ (Liu et al., 2024a) 11.27 ± 0.34 38.83 ± 0.17 24.03 ± 0.21 85.64 ± 0.32
+ECG-Byte 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
+5.3 Training Two-Stage Pretraining Methods in an End-to-End Fashion
+Although training methods are conventionally categorized as either two-stage or end-to-end, some works adopt an end-to-end strategy using an ECG-specific encoder architecture typically seen in two-stage methods (Wan et al., 2024).
+In Table 3, we simulate the approach of Wan et al.
+(2024) by jointly training the ECG encoder, learnable projection matrix, and LLM end-to-end using only an autoregressive objective.
+This contrasts with the two-stage method, where the ECG encoder is first trained with a self-supervised learning (SSL) objective and then frozen.
+We select strong ECG encoders from Table 1—namely, $L_{MIM}$ and $L_{MERL}$—and report performance on the ECG-QA MIMIC-IV and PTB-XL datasets (Oh et al., 2023).
+We denote the traditional 2-stage training approach for both ECG encoders as $L_{MIM}$ and $L_{MERL}$ and the end-to-end adaptation of 2-stage training approaches as End-to-End $L_{MIM}$ and End-to-End $L_{MERL}$ respectively in Table 3.
+The results clearly show a performance drop when the ECG encoder is not pretrained on a large dataset.
+This finding aligns with the Vision-Language Model (VLM) domain, where vision encoders are initially trained on internet-scale image data before integration with an LLM for visual understanding (Liu et al., 2023).
+Similarly, these findings indicate that ECG signals require extensive first-stage training with a dedicated self-supervised learning objective to learn robust representations before being applied to NLG.
+Even powerful, general-purpose pretrained encoders like ViT (Dosovitskiy et al., 2021) necessitate an initial training phase to effectively capture the unique characteristics of ECG data.
+In contrast, our method, ECG-Byte, compresses ECG signals using a rule-based algorithm (e.g., BPE) to represent them as tokens, enabling direct LLM training.
+Although these tokens belong to a different modality, we believe that the similarity in input representation allows the LLM to adapt more effectively without extensive first-stage training.
 5.4 Ablation Study
-We conduct several ablation studies to show the variability of performance with ECG-Byte when we alter the LLM used for finetuning, use different sequence lengths $L$ when inputting to the LLM, training ECG-Byte with various number of merges num_merges, and varying ECG lengths $T$.
-With the exception of the ablating parameter, we fix all other parameters to num_merges = 3500, $L=1024$, and $T=500$.
-We report results on the test set of the PTB-XL variant of ECG-QA (Oh et al., 2023) unless specified otherwise.
-[…]
-tab:diff_llm LLM BLEU-4 Rouge-L Meteor BertScore F1 GPT2 XL 1.5B (Radford et al., 2019) 12.30 ± 0.19 41.33 ± 0.57 26.48 ± 0.33 92.00 ± 0.06 Gemma 2B (Team et al., 2024) 13.78 ± 0.18 45.48 ± 0.55 28.32 ± 0.23 92.01 ± 0.02 OPT 1.3B (Zhang et al., 2022) 12.26 ± 0.20 41.84 ± 0.52 26.21 ± 0.29 91.78 ± 0.04 Llama 3.2 1B (Grattafiori et al., 2024) 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
+We conduct several ablation studies to show the variability of performance with ECG-Byte when we use varying LLMs during finetuning, sequence lengths $L$, and ECG lengths $T$.
+With the exception of the ablating parameter, we fix all other parameters to Llama 3.2 1B, num_merges = 3500, $L=1024$, and $T=500$.
+We report results when training and testing on the PTB-XL variant of ECG-QA (Oh et al., 2023) unless specified otherwise.
 Table 4: Ablation study on using different LLMs.
+LLM BLEU-4 Rouge-L Meteor BertScore F1
+GPT2 XL 1.5B (Radford et al., 2019) 12.30 ± 0.19 41.33 ± 0.57 26.48 ± 0.33 92.00 ± 0.06
+Gemma 2B (Team et al., 2024) 13.78 ± 0.18 45.48 ± 0.55 28.32 ± 0.23 92.01 ± 0.02
+OPT 1.3B (Zhang et al., 2022) 12.26 ± 0.20 41.84 ± 0.52 26.21 ± 0.29 91.78 ± 0.04
+Llama 3.2 1B (Grattafiori et al., 2024) 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
 Different LLMs
-We show the variability in performance of ECG-Byte when using different LLMs with similar numbers of parameters in Table LABEL:tab:diff_llm.
+We show the variability in performance of ECG-Byte when using different LLMs with similar numbers of parameters in Table 4.
 While Llama 3.2 1B (Grattafiori et al., 2024) achieves the best results, GPT2 XL 1.5B (Radford et al., 2019), Gemma 2B (Team et al., 2024), and OPT 1.3B (Zhang et al., 2022) also deliver comparable performances.
 These findings demonstrate that our method is not limited to Llama 3.2 1B but can achieve similar results across a variety of LLMs.
+Table 5: Ablation study on varying sequence lengths $L$.
+$L$ BLEU-4 Rouge-L Meteor BertScore F1
+512 13.61 ± 0.15 48.15 ± 0.57 29.10 ± 0.28 92.41 ± 0.05
+1024 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
+2048 13.88 ± 0.22 45.21 ± 0.48 28.31 ± 0.27 90.88 ± 0.02
+Table 6: Ablation study on varying lengths $T$.
+$T$ BLEU-4 Rouge-L Meteor BertScore F1
+250 12.64 ± 0.20 47.31 ± 0.26 27.97 ± 0.21 92.32 ± 0.06
+500 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
+1250 11.01 ± 0.19 43.84 ± 0.28 25.49 ± 0.20 93.07 ± 0.03
+2500 14.54 ± 0.17 48.03 ± 0.27 32.11 ± 0.22 92.91 ± 0.04
 Sequence Length
-Input lengths for LLMs are an important parameter to consider for efficient training since the calculation of attention is quadratic with respect to the input length (Vaswani et al., 2023).
-We present results on different sequence lengths $L$ in Table LABEL:tab:seq_len.
-Although the difference is not substantial, we can see that when $L=1024$ and $L=2048$ the model yields higher performance then $L=512$.
-We attribute this to the rate of truncation and padding for the encoded ECG.
-Observing Figure 2, most ECGs were being encoded to token sequence lengths of around 500 to 1500.
-Therefore, we hypothesize that when $L=512$ a large portion of the ECG token sequence gets truncated, resulting in lower performance.
+Input lengths for LLMs are critical for efficient training since the computation of attention scales quadratically with sequence length (Vaswani et al., 2023).
+Table 6 presents results for various sequence lengths $L$, which reveal minimal performance differences.
+As Figure 2 shows, most ECGs are encoded with token lengths around 500.
+Thus, we hypothesize that even small $L$ values (e.g., $L=512$) preserve the complete ECG information, resulting in only slight performance variations.
+We note that this can change depending on the initial ECG length $T$, in which we ablate in the next paragraph.
+We leave it up to future works for finding the sweet spot for the sequence length $L$ and its interaction with the initial ECG length $T$.
 […]
-tab:seq_len $L$ BLEU-4 Rouge-L Meteor BertScore F1 512 13.61 ± 0.15 48.15 ± 0.57 29.10 ± 0.28 92.41 ± 0.05 1024 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07 2048 13.88 ± 0.22 45.21 ± 0.48 28.31 ± 0.27 90.88 ± 0.02
-Table 5: Ablation study on varying sequence lengths $L$.
-Number of Merges
-The number of merges num_merges performed during training ECG-Byte corresponds to how much the algorithm compresses the concatenated sequence of quantized ECGS $\mathbf{X}_{\text{concat}}$.
-More num_merges means more compression, which can affect the expressiveness of the encoded sequence.
-In Table LABEL:tab:num_merges, we show the performance of our method with different num_merges.
-The results indicate that while performance varies slightly with the number of merges, values of num_merges greater than 500 generally yield similar outcomes.
-[…]
-tab:num_merges num_merges BLEU-4 Rouge-L Meteor BertScore F1 500 13.61 ± 0.53 46.50 ± 0.28 28.49 ± 0.49 92.33 ± 0.02 1750 14.50 ± 0.25 46.74 ± 0.48 30.03 ± 0.25 92.55 ± 0.01 2500 15.10 ± 0.39 46.37 ± 0.28 30.12 ± 0.23 92.53 ± 0.05 3500 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
-Table 6: Ablation study on varying number of merges num_merges.
 ECG length
-Lastly, we show the effect of the length $T$ being considered when encoding the ECG with ECG-Byte in Table LABEL:tab:len_ecg.
-We want to note that for the results of $T=2000$, the full unsegmented ECG is utilized.
+Lastly, we show the effect of the length $T$ being considered when encoding the ECG with ECG-Byte in Table 6.
+We want to note that for the results of $T=2500$, the full unsegmented ECG is utilized.
 Consequently, the number of instances available is less than the targeted dataset size of 400,000 (i.e., 97,244).
-Thus, when $T=2000$, we use the full dataset to train the model.
+Thus, when $T=2500$, we use the full dataset to train the model.
 For shorter segment lengths, such as $T=250$ and $T=500$, the model demonstrates strong performances indicating that shorter segments can effectively preserve relevant information for NLG.
-Interestingly, for $T=2000$, the model achieves the highest performance across all metrics.
+Interestingly, for $T=2500$, the model achieves the highest performance across all metrics.
 This suggests that when the model is trained with the full 10 second encoded ECG, it benefits from richer contextual information present in the complete ECG waveform.
-[…]
-tab:len_ecg $T$ BLEU-4 Rouge-L Meteor BertScore F1 250 12.64 ± 0.20 47.31 ± 0.26 27.97 ± 0.21 92.32 ± 0.06 500 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07 1250 11.01 ± 0.19 43.84 ± 0.28 25.49 ± 0.20 93.07 ± 0.03 2000 14.54 ± 0.17 48.03 ± 0.27 32.11 ± 0.22 92.91 ± 0.04
-Table 7: Ablation study on varying lengths $T$.
 5.5 ECG-Byte Analysis
@@
 Token Usage and Length Distribution
-[…]
 We examine the token usage and length distributions for ECG-Byte with num_merges = 3500 on a subsample of 277,840 ECGs from the PTB-XL dataset.
+For a 2-second ECG ($T=500$), the raw signal comprises $12\times 500=6000$ symbols.
+ECG-Byte then compresses these 6000 symbols at an average ratio of 12.66x, yielding approximately 500 tokens per ECG on average, as illustrated in Figure 2.
 The left panel of Figure 2 displays the token usage distribution, showing token frequency (y-axis) ranked in descending order (x-axis).
@@
 The right panel of Figure 2 illustrates the token length distribution of the encoded ECGs, with most falling between 500 and 1000 tokens, demonstrating ECG-Byte’s effective compression of the original signal.
-Additional examples of these distributions are provided in Appendix A.3.
+Additional examples of these distributions are provided in Appendix B.3.
+Figure 3: A mapping between tokens used for a given ECG Lead I. More examples are provided in Appendix B.2.
 Token to ECG Mapping
-Figure 3: A mapping between tokens used for a given ECG Lead I. More examples are provided in Appendix A.2.
 To illustrate how ECG-Byte encodes ECG signals, we analyze the mapping between tokens and signal features.
-Figure 3 shows an example Lead I ECG signal with unique token IDs (represented by different colors) overlayed.
+Figure 3 shows an example Lead I ECG signal with unique token IDs (represented by different colors) overlaid.
 The P wave, QRS complex, and T wave are distinctly captured by different tokens, though this precision varies across instances.
 As demonstrated, ECG-Byte effectively merges key regions of the signal.
-Additional examples are provided in Appendix A.2 due to page limitations.
+Additional examples are provided in Appendix B.2 due to page limitations.
+Figure 4: The attention weight overlaid on top of both the text (top) and ECG (bottom). More examples are provided in Appendix B.4.
 Attention Visualizations
@@
 We focus on one lead due to the uniformity of attention across encoded signal tokens.
-For interpretability, the reversed ECG signal is overlayed on the encoded ECG.
+For interpretability, the reversed ECG signal is overlaid on the encoded ECG.
 The model primarily attends to the textual portion of the input sequence, as shown in Figure 4.
@@
 We hypothesize that a similar phenomenon occurs in Figure 4, as the ECG tokens, though represented like text, are 1) newly introduced and 2) perceived as a different modality (e.g., vision).
-Additional examples are provided in Appendix A.4.
+We note that attention visualizations may not inherently indicate which parts of the input sequence contribute to the final generated output.
+Although methods such as LIME (Ribeiro et al., 2016) or Integrated Gradients (Sundararajan et al., 2017) may provide better attributions, we leave this investigation for future work.
+Additional examples are provided in Appendix B.4.
 6 Discussion and Conclusion
@@
 ECG-Byte introduces a paradigm shift in generative ECG language modeling by enabling efficient end-to-end training, compared to traditional two-stage pretraining approaches.
-Our pipeline demonstrates strong performances, achieving results comparable to two-stage methods while requiring only half the training time and approximately $48\%$ of the data.
+Our pipeline demonstrates strong performance, achieving results comparable to two-stage methods while being about 3 times faster in total training time and requiring approximately 48% of the data.
+Drawing inspiration from MEIT (Wan et al., 2024), we simulate two-stage pretraining for NLG in an end-to-end manner.
+Our findings indicate that the initial training stage is essential for achieving performance gains.
 In addition to its efficiency, ECG-Byte enhances interpretability.
-Figure 4: The attention weight overlayed on top of both the text (top) and ECG (bottom). More examples are provided in Appendix A.4.
 By analyzing its underlying mechanism, we observe that critical ECG regions, such as the P wave, the QRS complex, and the T wave, are effectively grouped during tokenization, as illustrated in Figure 3.
@@
 However, as shown in Figure 4, the model’s attention weight distribution resembles that of vision language models, focusing primarily on the textual components of the input sequence during generation.
-This work is in its early stages and needs further exploration.
-Future directions include: (1) refining BPE merging rules to better capture ECG-specific features, (2) adopting more advanced quantization techniques that preserve time-series characteristics (Carson et al., 2024b; Elsworth and Güttel, 2020; Carson et al., 2024a), (3) introducing stronger modality-specific distinctions, such as embeddings beyond [SIG_START] and [SIG_END] (Gui et al., 2023), and (4) extending ECG-Byte for conversational tasks through instruction tuning.
+We emphasize that the goal of ECG-Byte is not to claim it as the best method for representing ECG data or training generative, autoregressive Electrocardiogram-Language Models (ELMs).
+Instead, ECG-Byte demonstrates the potential of using a rule-based compressor rather than a learnable one for ECG data in the context of ELMs.
+As stated earlier, we find ECG-Byte both highly efficient and intuitive in compressing ECG data, and its tokenization is interpretable because it can be reversed to recover the original signal.
+We hope the broader research community will recognize the potential of rule-based compressors for generative tasks with different health signal data beyond ECGs.
+This early-stage work invites the community to contribute to advancing generative ECG language modeling by highlighting some future directions.
+Future directions include: (1) a comprehensive benchmark with different input representations and training methods for ELMs, (2) refining BPE merging rules to better capture ECG-specific features, (3) adopting more advanced quantization techniques that preserve time-series characteristics (Carson et al., 2024b; Elsworth and Güttel, 2020; Carson et al., 2024a), and (4) introducing stronger modality-specific distinctions, such as embeddings beyond [SIG_START] and [SIG_END] (Gui et al., 2023).
 Limitations
-One of the main limitations of this work is the scale in terms of computing and data.
-Since we only used a subset of the data to train and test our method, we were unable to train the model to its full potential.
-However, even with only using a small subset of the data, we are able to see extremely promising results compared to other 2-stage SSL pretraining methods.
-Therefore, we do not view this limitation as a major bottleneck.
+One key limitation of this work is the scale of both computing resources and data.
+Due to limited computing capacity, we trained on only a data subset using a batch size of 2 for a single epoch, which may have constrained the model’s full potential.
+Nevertheless, as prior work indicates that finetuned LLMs can perform satisfactorily with little data (Brown et al., 2020), the promising results of ECG-Byte—in comparison with other two-stage SSL pretraining methods—suggest that this limitation is not a critical bottleneck.
 […]
 This work is done in collaboration with the Mario Lemieux Center for Heart Rhythm Care at Allegheny General Hospital.
-We thank Wenhao Ding, Haohong Lin, Shiqi Liu, and Hyoeun Kang for the valuable discussions.
-Appendix A ECG-Byte
-A.1 Additional Pseudocode for ECG-Byte
-Input: Array of IDs ids, Pair to merge pair as $(u_{1},u_{2})$, New ID new_id
-Output: Merged vector of IDs
-Function merge(ids, pair, new_id):
-       Initialize empty vector new_ids with capacity of ids;
-       i $\leftarrow$ 0;
-       while i $<$ length of ids do
-[…]
-                   new_ids.append(new_id);
-                   i $\leftarrow\texttt{i}+2$;
-             end if
-            else
-                   new_ids.append(ids[i]);
-                   i $\leftarrow\texttt{i}+1$;
-             end if
-       end while
-      return new_ids;
-Algorithm 3 Merging a pair in an ID array
-Input: Array of IDs ids
-Output: HashMap of pairs and their frequencies
-Function get_stats(ids):
-       pair_counts $\leftarrow$ Parallel fold operation:;
-       foreach window of size 2 in ids do
-             Let $(u_{1},u_{2})\leftarrow$ elements of the window;
-             Increment the count of $(u_{1},u_{2})$ in local pair_count;
-       end foreach
-      pair_counts $\leftarrow$ Parallel reduce operation to combine local pair_count HashMaps;
-       return pair_counts;
-Algorithm 4 Calculating Frequency of Byte Pairs in an Array
-A.2 Mapping between Token and ECG
-We add more examples of the mapping between the ECG signal and the encoded tokens for ECG-Byte in Figures 5 and 6.
-Figure 5: A mapping between tokens used for a given ECG Leads I, II, III, aVL, aVR, aVF.
-Figure 6: A mapping between tokens used for a given ECG Leads V1, V2, V3, V4, V5, V6.
-A.3 Token usage and length distribution for varying num_merges
+We thank Wenhao Ding, Haohong Lin, Shiqi Liu, Hyoeun Kang, and Tony Chen for the valuable discussions.
+Appendix A Training Details
+A.1 Learning Rate Scheduler
+We utilize a custom learning rate scheduler in all of our experiments.
+This scheduler applies an initial learning rate $\mathrm{init\_lr}$ scaled by the model’s hidden dimension ($\mathrm{d_{model}}^{-0.5}$) and dynamically adjusts it based on training steps, with a warm-up phase of 500 steps.
+The learning rate at step $\mathrm{n_{steps}}$ is updated as $\text{lr}=\mathrm{init\_lr}\times\min\left(\mathrm{n_{steps}}^{-0.5},\ \mathrm{n_{warmup}}^{-1.5}\times\mathrm{n_{steps}}\right)$.
+Appendix B Additional Details Regarding ECG-Byte
+Figure 5: Plots of the token usage and length distributions for ECG-Byte where num_merges is 500, 1750, and 2500 from top to bottom.
+B.1 Additional Pseudocode for ECG-Byte
+We provide detailed pseudocode for the ECG-Byte encoding process, merge and get_stats functions in Algorithms 4, 2, and 3 respectively.
+B.2 Mapping between Token and ECG
+We add more examples of the mapping between the ECG signal and the encoded tokens for ECG-Byte in Figure 6.
+B.3 Token usage and length distribution for varying num_merges
 We add more examples of the token usage and length distributions for varying num_merges in Figure 5.
-Figure 7: Plots of the token usage and length distributions for ECG-Byte where num_merges is 500, 1750, and 2500 from top to bottom.
-A.4 Attention Visualizations
-We add more visualizations of the attention weights in Figures 8, 9, 10, 11, 12, 13, 14, 15.
-Figure 8: The attention weight overlayed on both text (top) and ECG (bottom).
-Figure 9: The attention weight overlayed on both text (top) and ECG (bottom).
-Figure 10: The attention weight overlayed on both text (top) and ECG (bottom).
-Figure 11: The attention weight overlayed on both text (top) and ECG (bottom).
-Figure 12: The attention weight overlayed on both text (top) and ECG (bottom).
-Figure 13: The attention weight overlayed on both text (top) and ECG (bottom).
-Figure 14: The attention weight overlayed on both text (top) and ECG (bottom).
-Figure 15: The attention weight overlayed on both text (top) and ECG (bottom).
-Appendix B 2-stage Pretraining Approaches
+B.4 Attention Visualizations
+We add more visualizations of the attention weights in Figure 7.
+Appendix C Two-stage Pretraining Approaches
 To be consistent, we normalize each ECG in the same manner as described in subsection 3.2.
@@
 Given these two features $I$ and $O$ we then describe the contrastive, masked, and dual approaches implemented for our baselines that are derived from commonly used techniques used throughout previous works (Oh et al., 2022; Choi et al., 2023; McKeen et al., 2024; Pham et al., 2024; Tang et al., 2024a, b; Vaid et al., 2022).
-B.1 Contrastive learning approaches
+C.1 Contrastive learning approaches
 We utilize a pretrained CLIP Radford et al.
@@
 This is formulated as:
-$\mathcal{L}_{\text{CL}}=-\frac{1}{N}\sum_{i=1}^{N}\Bigg{[}$ $\log\frac{\exp\left(\text{sim}(z_{i}^{\text{img}},z_{i}^{\text{txt}})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(z_{i}^{\text{img}},z_{j}^{\text{txt}})/\tau\right)}$
-$+$ $\log\frac{\exp\left(\text{sim}(z_{i}^{\text{txt}},z_{i}^{\text{img}})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(z_{i}^{\text{txt}},z_{j}^{\text{img}})/\tau\right)}\Bigg{]}$
+$\mathcal{L}_{\text{CL}}=-\frac{1}{N}\sum_{i=1}^{N}\Biggl{[}\log\frac{\exp\left(\text{sim}(z_{i}^{\text{img}},z_{i}^{\text{txt}})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(z_{i}^{\text{img}},z_{j}^{\text{txt}})/\tau\right)}+\log\frac{\exp\left(\text{sim}(z_{i}^{\text{txt}},z_{i}^{\text{img}})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(z_{i}^{\text{txt}},z_{j}^{\text{img}})/\tau\right)}\Biggr{]}$
 where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, and $\tau$ is a learnable temperature parameter.
@@
 […]
-B.2 Masked image modeling approaches
+C.2 Masked image modeling approaches
 Consider the normalized ECG image $I\in\mathbb{R}^{3\times C\times T}$ obtained as previously described.
@@
 […]
-B.3 Dual approaches
+Algorithm 2 Merging a Pair in an ID Array merge
+Input: Array of IDs ids, pair to merge pair as $(u_{1},u_{2})$, new ID new_id.
+Output: Updated array of IDs ids with merged pairs.
+1:  $i\leftarrow 0$, $write\leftarrow 0$
+2:  while $i<\text{len({ids})}$ do
+3:     if $i+1<\text{len({ids})}$ and $(\texttt{ids}[i],\texttt{ids}[i+1])=\texttt{pair}$ then
+[…]
+5:        $write\leftarrow write+1$
+6:        $i\leftarrow i+2$
+7:     else
+8:        Set $\texttt{ids}[write]\leftarrow\texttt{ids}[i]$
+9:        $write\leftarrow write+1$
+10:        $i\leftarrow i+1$
+11:     end if
+12:  end while
+13:  Truncate ids to length $write$
+14:  return ids
+C.3 Dual approaches
 The dual approach follows the previous two contrastive and masked image modeling approaches for pretraining the ECG encoder but simply just combines the losses like so:
@@
 […]
-Appendix C Additional Results
-C.1 Does Larger LLMs Yield Higher Performance?
-We present the results of ablating the size of the LLM in Table LABEL:tab:large_llm.
+Algorithm 3 Calculating Frequency of Byte Pairs in an Array get_stats
+Input: Array of IDs ids.
+Output: HashMap of byte pairs and their frequencies pair_counts.
+[…]
+2:     pair_counts $\leftarrow$ Empty HashMap
+3:     for all each window of size 2 in ids do
+4:        $(u_{1},u_{2})\leftarrow$ two elements of the window
+5:        Increment count for $(u_{1},u_{2})$ in pair_counts
+6:     end for
+7:  else
+8:     pair_counts $\leftarrow$ Parallel fold operation:
+9:     for all each window of size 2 in ids (in parallel) do
+10:        $(u_{1},u_{2})\leftarrow$ two elements of the window
+11:        Increment count for $(u_{1},u_{2})$ in the local HashMap
+12:     end for
+13:     Combine local HashMaps using a parallel reduce operation to obtain pair_counts
+14:  end if
+15:  return pair_counts
+Appendix D Additional Results and Discussions
+Table 7: Mean results with standard deviations over 5 random seeds on zero shot cross-dataset transferability.
+Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1
+$L_{CL}$ ECG-QA MIMIC-IV ECG-QA PTB-XL 11.64 ± 0.45 41.48 ± 0.11 25.74 ± 0.13 91.24 ± 0.05
+$L_{MIM}$ 11.70 ± 0.29 42.22 ± 0.28 26.41 ± 0.10 91.51 ± 0.03
+$L_{MERL}$ (Liu et al., 2024a) 11.53 ± 0.19 39.23 ± 0.40 25.58 ± 0.28 91.59 ± 0.03
+$L_{Dual}$ 9.71 ± 0.10 35.10 ± 0.28 24.91 ± 0.19 87.88 ± 0.08
+ECG-Byte 8.70 ± 0.04 40.39 ± 0.40 23.29 ± 0.18 91.51 ± 0.03
+$L_{CL}$ ECG-QA PTB-XL ECG-QA MIMIC-IV 5.10 ± 0.04 22.77 ± 0.28 14.63 ± 0.32 77.89 ± 0.13
+$L_{MIM}$ 7.68 ± 0.46 35.77 ± 0.13 22.32 ± 0.33 90.28 ± 0.07
+$L_{MERL}$ (Liu et al., 2024a) 7.39 ± 0.15 28.33 ± 0.58 18.59 ± 0.35 89.30 ± 0.05
+$L_{Dual}$ 7.49 ± 0.21 30.53 ± 0.59 20.25 ± 0.27 86.53 ± 0.11
+ECG-Byte 7.86 ± 0.13 35.01 ± 0.41 21.49 ± 0.24 90.29 ± 0.07
+D.1 Cross Dataset Transferability
+We present the results of cross-dataset transferability in Table 7, comparing our approach, ECG-Byte, with two-stage pretraining methods.
+ECG-Byte achieves the best zero-shot transfer performance in BLEU-4 and BertScore F1 scores when transferring from the ECG-QA PTB-XL dataset to the ECG-QA MIMIC-IV dataset.
+When transferring from the ECG-QA MIMIC-IV dataset to the ECG-QA PTB-XL dataset, although other two-stage pretraining methods demonstrate higher performance, ECG-Byte maintains competitive results across all metrics.
+Table 8: Ablation study on varying number of merges num_merges.
+num_merges BLEU-4 Rouge-L Meteor BertScore F1
+500 13.61 ± 0.53 46.50 ± 0.28 28.49 ± 0.49 92.33 ± 0.02
+1750 14.50 ± 0.25 46.74 ± 0.48 30.03 ± 0.25 92.55 ± 0.01
+2500 15.10 ± 0.39 46.37 ± 0.28 30.12 ± 0.23 92.53 ± 0.05
+3500 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
+Number of Merges
+The number of merges num_merges performed during training ECG-Byte corresponds to how much the algorithm compresses the concatenated sequence of quantized ECGS $\mathbf{X}_{\text{concat}}$.
+More num_merges means more compression, which can affect the expressiveness of the encoded sequence.
+In Table 8, we show the performance of our method with different num_merges.
+The results suggest that, although performance fluctuates slightly with the number of merges, varying num_merges generally produces comparable outcomes.
+D.2 Does Larger LLMs Yield Higher Performance?
+We present the results of ablating the size of the LLM in Table 9.
 Interestingly, the performance across the three different model sizes (1B, 3B, 8B) remains fairly similar.
 We believe that the limited dataset size prevents the larger models from realizing their full performance potential.
-We hypothesize that increasing the amount of training data would enable the larger models to leverage their greater capacity, resulting in observable performance improvements.
+We hypothesize that increasing the amount of training data would enable the larger models to leverage their greater capacity.
+Table 9: Ablation study on how larger LLMs perform for NLG.
+LLM BLEU-4 Rouge-L Meteor BertScore F1
+Llama 3.2 1B (Grattafiori et al., 2024) 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
+Llama 3.2 3B (Grattafiori et al., 2024) 14.80 ± 0.17 46.55 ± 0.21 29.53 ± 0.16 92.42 ± 0.01
+Llama 3.1 8B (Grattafiori et al., 2024) 13.80 ± 0.16 46.29 ± 0.25 28.56 ± 0.11 92.44 ± 0.05
+D.3 Encoder-Free Vision-Language Models
+Released in 2023, Fuyu-8B (Bavishi et al., 2023) was a seminal work that introduced the concept of encoder-free Vision-Language Models (VLMs).
+To align the image modality with text, Fuyu-8B first patches an image and then projects these patches linearly into the first layer of a decoder-only transformer network (Vaswani et al., 2023).
+Fuyu-8B demonstrated strong performance on visual understanding tasks, and since its release, several works have explored encoder-free architectures for VLMs (Diao et al., 2024; Wang et al., 2025).
+However, at the time of this study, most VLMs continue to rely on vision encoders, so we did not pursue additional experiments applying encoder-free methods to ECGs.
+Nonetheless, there is a clear parallel in motivation between encoder-free VLMs and ECG-Byte.
+Utilizing ECG-specific encoders—whether explicitly designed for ECGs or pretrained on internet-scale data—can introduce strong inductive biases into the learned representations (Diao et al., 2024).
+In the encoder-free VLM literature, simple and low-overhead neural networks such as linear projections (Bavishi et al., 2023), MLPs (Wang et al., 2025), and convolutional networks (Diao et al., 2024) are used to align visual inputs with language models.
+In contrast, ECG-Byte employs a purely rule-based method—Byte Pair Encoding (BPE) (Gage, 1994)—to tokenize ECGs directly, enabling seamless integration with language model input alongside text tokens.
+While learning-based compressors with low overhead as seen in the encoder-free VLM literature may hold potential, we observe that even without any learnable modules to encode the ECG signal, ECG-Byte achieves strong performance.
+Therefore, we leave the exploration of incorporating encoder-free VLM concepts into the development of Electrocardiogram-Language Models (ELMs) to future work.
+D.4 Qualitative NLG Examples
+We provide qualitative NLG examples of successful (Figure 9) and unsuccessful generations (Figure 8).
+Algorithm 4 Encoding Process for ECG-Byte
+Input: Input $X_{\text{symb}}$ and merge history merges.
+Output: Vector of encoded IDs output_ids.
+1:  ids $\leftarrow$ Convert $X_{\text{symb}}$ to a vector of u32
+2:  trie_root $\leftarrow$ Initialize root TrieNode
+3:  for $b\leftarrow 0$ to $255$ do
+4:     insert(trie_root, [b], b)
+5:  end for
 […]
-tab:large_llm LLM BLEU-4 Rouge-L Meteor BertScore F1 Llama 3.2 1B (Grattafiori et al., 2024) 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07 Llama 3.2 3B (Grattafiori et al., 2024) 14.80 ± 0.17 46.55 ± 0.21 29.53 ± 0.16 92.42 ± 0.01 Llama 3.1 8B (Grattafiori et al., 2024) 13.80 ± 0.16 46.29 ± 0.25 28.56 ± 0.11 92.44 ± 0.05
-Table 8: Ablation study on how larger LLMs perform for NLG.
-C.2 Qualitative NLG Examples
-We provide qualitative NLG examples of successful (Figure 17) and unsuccessful generations (Figure 16).
+7:     insert(trie_root, token_sequence, token_id)
+8:  end for
+9:  output_ids $\leftarrow$ Empty list
+10:  $i\leftarrow 0$
+11:  while $i<\text{len({ids})}$ do
+[…]
+[…]
+[…]
+[…]
+16:        $\texttt{id}\leftarrow\texttt{ids}[j]$
+17:        if id exists in node.children then
+[…]
+[…]
+[…]
+[…]
+22:           end if
+23:        else
+24:           break
+25:        end if
+26:     end for
+[…]
+28:        Append match_id to output_ids
+[…]
+30:     else
+31:        Append $\texttt{ids}[i]$ to output_ids
+32:        $i\leftarrow i+1$
+33:     end if
+34:  end while
+35:  return output_ids
+Figure 6: A mapping between tokens used for a given ECG Leads I, II, III, aVL, aVR, aVF, V1, V2, V3, V4, V5, V6..
+Figure 7: The attention weight overlaid on both ECG (left) and text (right).
 Figure 8: Randomly sampled NLG results of unsuccessful generations on the PTB-XL test set from ECG-QA.